### v1

# Towards Fully Autonomous Research Powered by LLMs: Case Study on Simulations

Abstract
The advent of Large Language Models (LLMs) has created new opportunities for the automation of scientific research, spanning both experimental processes and computational simulations. This study explores the feasibility of constructing an autonomous simulation agent (ASA) powered by LLM, through sophisticated API integration, to automate the entire research process, from experimental design, remote upload and simulation execution, data analysis, to report compilation. Using a simulation problem of polymer chain conformations as a case study, we assessed the performance of ASAs powered by different LLMs including GPT-4-Turbo. Our findings revealed that ASA-GPT-4o achieved near-flawless execution on designated research missions, underscoring the potential of LLMs to manage complete scientific investigations autonomously. The outlined automation can be iteratively performed up to twenty cycles without human intervention, illustrating the potential of LLMs for large-scale autonomous research endeavors. Additionally, we discussed the intrinsic traits of ASAs in managing extensive tasks, focusing on self-validation mechanisms and the balance between local attention and global oversight.
Introduction
In recent years, the rapid development of artificial intelligence technology, particularly the emergence of large language models (LLMs), has brought revolutionary changes to scientific research[1, 2, 3, 4, 5]. Since the launch of ChatGPT by OpenAI, LLMs such as GPT-4 and Anthropic’s Claude have demonstrated outstanding performance in natural language processing, text generation, and translation [6, 7]. These AI models not only excel in traditional text processing tasks but also show significant potential in scientific research, gradually transforming conventional research methods and providing scientists with unprecedented tools and methodologies[8, 9, 10, 11].
The application of artificial intelligence (AI) in the fields of chemistry and materials engineering is continually expanding. By integrating machine learning and automated experimental technologies, faster experimental design and data analysis can be achieved, leading to the discovery of new materials and the optimization of existing material properties in a shorter timeframe [12, 13, 14]. In the domain of physical chemistry knowledge parsing, Meta’s Galactica assists researchers with tasks such as information organization, knowledge inference, and writing, showcasing the immense potential of large language models in scientific and professional domains [13]. Meanwhile, DeepMind’s AlphaFold series of algorithms enable the prediction of structures for biological molecules, from proteins and nucleic acids to larger and more complex complexes [15, 16]. In the fields of chemical synthesis and drug design, AI aids in complex chemical synthesis planning, chemical property prediction, and the generation of new materials. Integrating LLMs with specialized chemical tools and databases can extend their capabilities beyond general language tasks [17, 18, 19, 20, 21]. Additionally, some studies have highlighted the crucial role of prompt engineering in guiding models to produce accurate and useful outputs. Techniques such as Chain of Thought, ReAct, and Tree of Thoughts help structure the model’s reasoning process, reduce errors, and improve the quality of generated results [20, 22, 23, 24].
Recent research shows that continuous human-AI interaction can accelerate various stages of scientific research. However, this human-led, AI-assisted model still limits the speed and depth of scientific research. Some work has explored an AI-dominated research model, such as smart laboratories (SDLs) that integrate AI, laboratory automation, and robotics to conduct autonomous experiments. These systems can automatically adjust and optimize based on intelligent analysis and decision-making support from experimental results [25, 26, 27]. The ultimate goal for future AI scientists is to achieve full automation of scientific research processes, from experimental design to data analysis and conclusion generation. This end-to-end automation is set to become a major development in chemical research and a primary objective in chemical engineering, including both experimental and simulation aspects.
This paper explores the AI automation paradigm in polymer theory simulation research. In traditional simulation research, a supervisor typically discusses and plans the research project with students. The students then write and modify the simulation program based on physical models, run simulations under different parameter conditions on remote servers, collect data, analyze results, and write reports. This paper envisions a new research paradigm where, after the research has been planned, the subsequent tasks are solely carried out by AI. Using an autonomous simulation agent (ASA) powered by LLM, after human researchers input a research plan (RP), the ASA will then take on the role of the student to complete the entire research process. Humans only need to review the AI’s results at the end, thus achieving a research mode without human intervention.
To test this research paradigm, we systematically attempted ASAs powered by different LLMs on a moderately complex simulation problem in polymer physics. This involved an AI fully autonomously coding the simulation program, remote uploading and executing simulations, data organization, plotting, and scientific report writing.
Random walk and self-avoiding walk models are of significant physical importance in describing the spatial configuration of polymer chains. The random walk model assumes that each ‘step’ of the chain is independent and random, while the self-avoiding walk model considers the volume exclusion effects between chain segments. Therefore, they exhibit different dynamic behaviors and scaling properties, with different scaling relationships between the mean square end-to-end distance and the number of chain segmentss N. This scaling relationship has been theoretically derived by Flory, De Gennes, and others [28, 29] and validated through simulations and experiment [30, 31]. This problem is fundamental in polymer physics and has been extensively studied. However, for AI, it still presents a novel challenge. The difficulty of predicting the scaling exponent makes it neither too easy nor prohibitively difficult, thus serving as an excellent model problem for testing AI capabilities. We prompted the AI to write simulation programs for both random walk and self-avoiding walk models, compute the mean square end-to-end distance for different chain segment numbers N, fit the results to $\left\langle R^{2}\right\rangle\propto N^{\nu}$ , determine the scaling exponent $\nu$, plot the data, and write a complete scientific report.
The ASA leverages an API automation program (AutoProg) to enable multi-round text interactions with LLMs via API (Figure 1). Initially, the AutoProg provides an RP witten by human researchers to an LLM, and the LLM will break the plan down into multiple sub-tasks and write Python code to complete the first sub-task in its response. The AutoProg then extracts the Python code, executes it, and returns the execution results and any errors along with the conversation history back to the LLM. The LLM then decide whether to modify the Python code or to proceed to the following sub-tasks. This process repeats until the LLM determines that the mission is complete, at which point the AutoProg stops. Upon mission completion, humans evaluate its success based on criteria such as correct simulation, plotting, etc. The ASA we developed is a general system (minor adjustments may be required in the AutoProg for different LLM APIs) that requires only a change in the original RP to be applicable to any simulation problem involving Python programming. In addition to the RPs designed for polymer conformation simulations, we have also crafted RPs for solar system planetary orbit simulations and asteroid trajectory simulations (Figure 1C and SI).
Figure 1: A. Schematic diagram of the AI automation process for theoretical simulation research conducted in this study. The human researcher writes the research plan (RP) for the theoretical simulation and provides it to the autonomous simulation agent (ASA), which is developed by sophisticated API programming (see Method section). The ASA first generates the simulation program, then uploads it to a remote server, performs simulation calculations with different parameter conditions, collects data, and finally writes a report based on the simulation results. B. Diagram of the LLM dialogue. The human researcher only inputs the RP at the beginning, and then the ASA makes decisions based on the dialogue history, achieving a fully autonomous research mode. C. Two typical simulation problems studied in this paper: the chain conformation simulation of a random walk and the gravitational model simulation of asteroid orbit prediction (See SI).
We compared the performance of ASAs powered by various LLMs, including GPT-4o, and explored common characteristics and issues when LLMs tackle long tasks. These findings provide insights for researchers to better improve and utilize LLMs for AI-driven scientific research.
Overall, this paper presents the first fully autonomous AI-conducted simulation research process powered by LLMs. By adopting this new research paradigm, we aim to explore the potential of AI in highly complex and data-intensive scientific research, breaking traditional human limitations and advancing the automation and intelligence of scientific research.
Figure 2: Overview of RP 1-3. We designed three RPs, each containing multiple steps such as simulation, plotting, and report writing, and provided them to the ASA. RP 1 and RP 3 required the ASA to generate a complete Python program to simulate polymer chains, run the simulation locally, and produce a final report. RP 2 required the ASA to modify the provided simulation program, run the simulation on a remote server, and then generate a report. The provided simulation program (omitted in the figure) and remote server information were included in RP 2.
Results
Automating Basic Simulations in Research
We designed three simulation missions related to polymer chain conformation modeling to demonstrate and test the ability of ASAs powered by different LLMs to autonomously complete the entire research process. Each mission includes multiple sub-tasks, covering the full spectrum from conducting experiments to analyzing data and writing reports. The research plans (RPs) for these missions are shown in Figure 2. Among these RPs, RP 1 is the simplest involving only basic research steps for simulations given below.
  • RP 1 requires generating a Python program to simulate a random walk, sampling different numbers of chain segments N, deriving the scaling relation $\left\langle R^{2}\right\rangle\propto N^{v}$, saving chain conformation graphs and scaling relation fit plots, and writing a research report.
We established several evaluation criteria based on the requirements of the RP, such as whether graphs and plots were generated and whether the derived scaling relationships were correct, to assess task completion. We tested nine ASAs powered by different LLMs, including GPT-4o, conducting 20 trials for each RP and recording the achievement of the evaluation criteria across these trials, as shown in Figure 3.
ASA-GPT-4o and ASA-GPT-4-Turbo excelled on RP 1, achieving near-perfect completion in both simulation and report writing tasks (Figure 3B). ASA-Claude-3.5 and ASA-Qwen-2 followed, performing well in simulation tasks but sometimes generating reports in response or as text files, failing to meet RP 1’s requirements. Other models performed well in plotting graphs but lacked accuracy in establishing correct physical models and deriving correct scaling values through simulations.
Although chain conformation simulations are convenient for comparing different LLMs, they are less convincing due to being previously well-studied topics. To make a stronger case, we need more challenging problems. Additionally, to demonstrate that ASA can solve problems across different scientific domains, we designed RPs for gravitational simulation problems corresponding to RP S1 and RP S2 (Figure S1 and Figure 1C). The mission of RP S2 is to evaluate the risk posed by a meteorite (or asteroid) to earth, given its initial position and velocity and write a report. This mission is challenging even for graduates from physics departments. By simply inputting RP S2 into the same ASA used for chain conformation problems, the ASA will automatically acquire the necessary parameters of the solar system, code the simulation program to calculate the asteroid’s trajectory, and ultimately write the report. This result underscores the effectiveness of our API programming strategy (see Method section) and strongly demonstrates that ASA can address problems in entirely different fields.
Automating Simulations with Remote Server Operations and Volume Exclusion
RP 2 increases complexity with server interactions and RP 3 doubles the mission length and simulation difficulty. Brief introductions of these two RPs are shown below (Figure 2).
  • RP 2 directly provides a random walk simulation program, asking the ASA to modify it, run simulations in a designated folder on a remote server, download the experimental data, and generate graphs and plots and a research report.
  • RP 3 is similar to RP 1 but includes both random walk and self-avoiding walk simulations.
For RP 2, which involves remote upload and remote simulations, completion rates across ASAs significantly declined, with some agents consistently failing across multiple trials. ASA-Claude-3.5 achieved the highest completion rate, successfully completing the entire mission, including generating all required graphs in over two-thirds of trials, though it occasionally missed displaying some generated graphs in the report. ASA-GPT-4o, ASA-GPT-4-Turbo, and ASA-Qwen-2 also demonstrated some success in solving complex missions over 20 trials (Figure 3).
RP 3 required simulating a self-avoiding walk based on the random walk simulation in RP 1. Although the process was longer, it was less complex than RP 2, resulting in improved performance across agents. ASA-GPT-4o, ASA-GPT-4-Turbo, ASA-Claude-3.5, and ASA-Gemini-1.5-Pro performed relatively well. Notably, none of the agents provided a correct self-avoiding scaling due to incorrect sampling methods during simulation, which will be discussed in detail in the Discussion section.
Moreover, ASA-GPT-4o and ASA-Claude-3.5 demonstrated unique advantages in generating rich content. They frequently produced detailed reports exceeding 1,000 words, delving into the distinctions between RW and SA scaling, and occasionally included precise literature citations. ASA-Claude-3.5 referenced accurate self-avoiding walk scaling relationships validated through theoretical derivations and experimental verifications, accompanied by brief analyses of simulation result deviations. Additionally, ASA-GPT-4o presented experimental data in enriched formats, such as frequency distribution plots of end-to-end distances and interactive web-based charts.
Automating Agent Coordination through RP Generation
In the preceding section, RPs were provided by humans. In this section, we task a Main AI with automatically breaking down a human-provided RP into sub-tasks and distributing them as AI RPs to several Subordinate AIs.
We provided RP 1 and 2 directly to the Master AI, adjusting the ASA to ensure each sub-task was handled by a distinct Subordinate AI without access to the Main AI’s or other Subordinate AIs’ conversation histories. The Main AI solely presents the AI RPs (Figure 4A) and received reports upon task completion.
We also employed GPT-4-turbo for this experiment. The results (SI-data-2) revealed that for RP 1, the Main AI effectively decomposes the mission into sub-tasks and successfully distributed them to Subordinate AIs, thereby completing RP 1. However, for RP 2, the Main AI encounters repeated failures. The primary issue lies in precise information transmission. RP 2 provides a random walk simulation program and details of a remote server, including hostname and username. Despite multiple attempts, the Main AI struggles to convey both the simulation program and server details accurately to the Subordinate AIs (Figure 4B). We attempted to instruct the Main AI to include all relevant information in the AI RPs, but this did not notably enhance the accuracy of information transmission. Additionally, we observed instances where AI RPs contained irrelevant information, potentially confusing the Subordinate AIs about the mission’s focus.
Automating the Automation: Crafting Multi-tier RPs for Large-Scale Research
Back in the Automating Basic Simulations section, for each ASA and each RP, we undertake 20 rounds of ‘automation’ by executing RP 1 and collecting the results including codes and files generated by the ASA. Can we automate the entire process of this study on reaserch automation or can we even automate the automation itself?
RP 4 is designed for this purpose. It is nested within RP 1, instructing the Primary AI to execute the command “python AutoProg.py -s p1.txt -n i” 20 times. Each execution generates an Agent AI to fulfill RP 1 contained in p1.txt, while the Primary AI collects all generated files and conducts result analysis. The detailed content of RP 4 is depicted in Figure 4C.
Using ASA-GPT-4-Turbo, we conducted tests on RP 4. The results (SI) strikingly demonstrate that the Primary AI successfully executed the 20 instances of RP 1, meticulously organizing all files into their respective folders and providing a brief summary after all experiments. During this process, the AI independently and without interruption wrote 66 programs (including 26 error versions), conducted more than 20 simulations, generated 120 images, and wrote 20 research reports totaling over 5,000 words. Marvelously, all these missions were accomplished without any human intervention. This illustrates the remarkable capability of multi-tier RP structures in handling large-scale missions.
Figure 3: Completion Rates of Various ASAs for RP 1-3. A. Table of LLM models used in this study. B. Statistics of the number of criteria met by each ASA over 20 experiments. We established seven criteria for RP 1-3 to measure mission completion rates and counted the number of times each ASA met these criteria across 20 experiments. Some zero results are not displayed. C. ASA Scoring. Relative scores for each ASA in RP 1-3 were calculated using the EWM and TOPSIS methods. D. ASA Generated Research Report Example. Demonstrates a Word report generated by ASA-GPT-4 for RP 1, including four sections, chain conformation diagrams and a scaling relation fit plot (also see SI).
Figure 4: A. Example of Main AI’s Response. The Main AI first breaks down the overall mission into sub-tasks, then generates AI RPs beginning with ’< < < prompt’ and ending with ’end > > >’. The AI-generated RPs are then passed to subordinate AIs for execution. B. Example of Faulty and Incomplete AI RP. The black box highlights where the AI RP fails to provide necessary information such as username and password for connecting to the remote server, and the red box highlights where the AI RP includes irrelevant information. C. Overview of RP 4. The command ’python AI4SCI _VE13.py -s p1.txt -n i’ generates an Agent AI to complete the mission of RP 1, where AI4SCI _VE13.py is the API automation execution program, p1.txt is the text file containing the RP, and i denotes the $i^{th}$ execution of the command.
Discussion
The experiments demonstrate that LLMs, with their robust coding ability [24, 32, 33, 34], can be directly employed in developing agents to autonomously complete complex research missions when aided by proper API programming. ASA-GPT-4o and ASA-GPT-4-turbo, notably, exhibits near-perfect completion rates for simpler missions, highlighting their potential. Multi-tier RP designs effectively handle more complex missions, though accurate information transmission remains challenging for such RPs. Enhancements in RP design and AI coordination are essential for fully harnessing AI’s potential in automating scientific research. Below, we discuss several problems existing in the current reaserch.
Common Issues in Sequential Task
Despite varying completion rates among tested ASAs, their errors and failure points share common characteristics, reflecting the inherent logical features of LLMs.
First, some ASAs tend to overlook parts of the RP’s requirements or “take shortcuts”, not fully following the RP’s instructions. For instance, they might skip plotting, fail to write a report, or ask humans to execute part of the task. These issues affect mission completeness and can disrupt the entire workflow. To improve autonomy in AI for science, we need to address these issues by continuously generating outputs describing the current mission status, periodically reviewing generated data, and confirming strict adherence to mission requirements. These improvements can be integrated through AutoProg modifications or by incorporating them into the "system" prompt input to LLMs or fine-tuning the LLMs.
Second, ASAs often underperform on technical aspects of tasks due to a lack of specialized domain knowledge. For example, when simulating a random walk, the algorithm needs to generate unit vectors uniformly distributed on a sphere to ensure isotropy. Although ASAs can often generate spatially random unit vectors, they fail to ensure uniform distribution on the sphere. In another example, for simulating self-avoiding walks, ASAs often incorrectly extend the random walk strategy by merely checking distances between points, resulting in high-energy state samples with large statistical errors. Correct sampling should use importance sampling to generate statistically significant conformations[35]. To avoid these issues in research, reliable solutions should be provided in the RPs, or relevant domain knowledge should be used for fine-tuning the LLMs to enhance their understanding.
Error Loops in Debugging
Each AutoProg involves a Python code debugging process. When bugs are detected, the LLM receives error messages and attempts to revise the code, iterating until it executes correctly. However, we’ve noted instances where the LLM struggles to generate functional code, resulting in mission failures. This typically occurs when the LLM becomes “stuck” after repeated unsuccessful debugging attempts, cycling through the same erroneous code without progress.
This phenomenon may be associated with the LLM’s propensity to depend heavily on past dialogue content. Research has documented instances where LLMs utilize context to deliver responses preferred by users, indicating an excessive reliance on previous information [36, 37, 38]. While this capability allows LLMs to adapt responses based on short dialogue histories, it can also lead to persistent errors.
In our experiments, to prevent such deadlocks, we implemented a maximum debug attempt limit. Once this limit is reached, the AutoProg returns a “mission failed” message. To ensure the smooth execution of concrete research missions, it is crucial to avoid these debugging loops. One solution involves clearing parts of the memory after a certain number of debug attempts or encouraging the LLM to generate varied content in each iteration (Some LLMs include parameters to avoid generating duplicate content and to adjust the creativity of the models.).
Balancing Global Oversight and Local Attention
In our experiments, ASAs exhibited a significant level of global oversight, which allows them to track mission progress based on the dialogue history. They can retain details from earlier stages of missions even after focusing on lengthy sub-tasks. However, as discussed above, ASAs occasionally skip steps in prolonged missions or fail to deliver comprehensive solutions for sub-tasks, indicating a lack of sufficient local attention.
Achieving successful and precise mission completion requires a delicate balance between global oversight and local attention, two perspectives that are normally conflicting.
In Automating Agent Coordination section, we designed a collaboration plan where a Main AI and some Subordinate AIs handled different aspects of the mission: the Main AI defined sub-task requirements and received reports, while Subordinate AIs focused on specific sub-tasks. This approach minimized redundant information in their dialogue history, enhancing the Main AI’s global awareness and the Subordinate AIs’ local attention. While the Main AI effectively completed the mission in RP 1, the Main-Sub AI model encountered challenges in RP 2 due to the failure of precise information transmission. One potential optimization is to involve human-designed RPs distributed among various Subordinate AIs for execution.
Limitations and Next Steps
This study demonstrates the capability of LLM powered research agent ASA to independently undertake sophisticated simulation research missions, using a well-recognized polymer physics simulation problem and a more challenging celestial-simulation problem as representative examples. However, there remains ample opportunity to explore ASA’s potential in tackling novel scientific challenges. Additionally, this study focuses on presenting an automated workflow, leaving room for further optimization of prompts and automation programs.
Several inherent limitations of ASAs were identified: (1) ASAs may overlook mission requirements, necessitating methods to prioritize critical information; (2) general LLMs lack specialized domain knowledge, requiring supplementary information in RPs; (3) LLMs may persist in errors by relying too heavily on previous content, requiring memory management strategies in ASAs’ development; (4) balancing global awareness and local attention poses a challenge; (5) accurately transmitting specific information remains problematic.
Addressing these issues is crucial for enhancing the role of AI in scientific research. This requires a thorough analysis of these limitations and adjustments to research methodologies and LLM capabilities, leveraging their strengths while mitigating their weaknesses.
Conclusion
This study demonstrates the potential of LLM powered research agent ASA to autonomously conduct scientific research missions, using a polymer theory simulation problem and a more challenging asteroid-trajectory prediction problem as model problems. This process, encompassing simulation coding, remote execution, data analysis, and report generation, represents the first fully autonomous AI-directed simulation research. The notable success herein can be largely attributed to the sophisticated API programming. Agents such as ASA-GPT-4o demonstrate high completion rates for missions, a strong grasp of task requirements, and consistent alignment with overall objectives across extensive tasks. Notably, the performance exhibited in completing RP 4, entailing a hierarchy of nested tasks, surpassed expectations in a significant manner. These capabilities underscore the readiness of ASAs to independently manage long-term scientific research missions, offering a new perspective on AI applications in science as well as in chemical engineering.
Method
API automation Program (AutoProg)
Based on the official documentation of various LLMs, we implemented the API automation programs (AutoProg) by configuring Python API calling methods and setting up the local environment. This setup included installing necessary Python packages like numpy, matplotlib, paramiko, and python-docx tailored for our simulation tasks. An AutoProg serves dual functions: 1) it engages in dialogue with the LLM via the API and preserves all historical dialogues, and 2) it extracts and executes Python programs generated by the LLM in response to prompts.
For instance, in the case of RP 1, an AutoProg initiates the automated research process by reading mission requirements saved in a txt file. It then calls the API to transmit the RP to the LLM, instructing it to generate a comprehensive Python program. The AutoProg extracts this program from the LLM’s response, sends it back to the LLM via the API along with entire dialogue history, and requests error checking and compliance verification against mission requirements. Subsequently, the AutoProg executes the validated Python program, capturing and storing output or errors. In case of errors, the AutoProg sends the prior dialogues and program errors back to the LLM, asking for modifications. This process repeats until the program runs correctly. If errors persist beyond a predefined threshold, the AutoProg halts with a “mission failed” output. Conversely, upon successful execution, it returns the program output to the LLM, prompting it to check the dialogues and files in the working directory to determine the task progress. If the task remains incomplete, the next Python program will be generated by the LLM to continue the mission, iterating through checking and debugging processes until completion. When the mission achieves full execution, the AutoProg concludes with a "mission complete" output.
AutoProg.py and RP.txt, containing predefined instructions, are stored in the local working directory. To initiate the automated research process for each experiment, users navigate to the directory via the command line (CMD) and execute commands such as “python AutoProg.py -s RP.txt”. Alternatively, AutoProg can be elevated to a command by creating a batch file (Windows) or shell script (Linux), allowing users to simply enter “AutoProg RP.txt” from any directory to execute it. A detailed video recording of the execution process is available for reference (SI-video), in which a batch file is used to facilitate the execution by running the command “do AI4S_prj1.txt”.
To evaluate LLM performance across RP 1-3, each ASA underwent 20 tests per RP, with all generated files and AutoProg outputs saved. Sample results are provided in the data (SI-data-1).
ASA Scoring via EWM and TOPSIS
To evaluate the performance of ASAs on designated RPs, we devised specific criteria for RP 1-3, including diagram generation, report composition, etc. Through systematic observation across multiple trials, we recorded each agent’s fulfillment against these criteria. For quantitative analysis and scoring, we deployed the Entropy Weight Method (EWM) and TOPSIS (Technique for Order of Preference by Similarity to Ideal Solution).
The EWM, a well-established approach for determining the relative significance of criteria in decision-making [39], was utilized to assign weights based on data variability. This ensures criteria with higher variability, and thus more informational value, receive higher weights in the final evaluation.
TOPSIS, originally formulated by Hwang and Yoon [40], was subsequently applied. This method computes the geometric distance between each agent’s performance and both ideal and anti-ideal solutions, facilitating a comparative assessment of ASA efficacy.
The evaluation procedure is delineated as follows.
Initially, fulfillment data for all ASAs across various criteria were gathered over 20 trials, encapsulated in matrix X, wherein $x_{ij}$ signifies the frequency of criterion j’s attainment by agent i. Normalization was then performed on $x_{ij}$:
$r_{ij}=\frac{x_{ij}-min(x_{ij})}{max(x_{ij})-min(x_{ij})},$ (1)
yielding the normalized matrix $\mathbf{R}=[r_{ij}]$. Subsequently, criterion weights were established utilizing EWM. The relative proportion $p_{ij}$ of the criterion j:
$p_{ij}=\frac{r_{ij}}{\sum_{i=1}^{m}r_{ij}}.$ (2)
The entropy $e_{j}$ of the criterion j:
$e_{j}=-k\sum_{i=1}^{m}p_{ij}\ln(p_{ij}),$ (3)
with $k=1/\ln(m)$, and $m$ denoting the total number of agents.
The weight $\omega_{j}$ of the criterion $j$:
$\omega_{j}=\frac{d_{j}}{\sum_{j=1}^{n}d_{j}},$ (4)
with $d_{j}=1-e_{j}$ the disparity of the criterion and $n$ the total number of criteria. Conclusively, agent scores were derived via TOPSIS. Multiplication of $r_{ij}$ by its respective weight $\omega_{j}$ yields the weighted normalized matrix $\mathbf{V}=[v_{ij}]$ with $v_{ij}=\omega_{j}\cdot r_{ij}$. The positive ideal solution $v_{j}^{+}$ and negative ideal solution $v_{j}^{-}$ were determined as $v_{j}^{+}=max(v_{ij})$ and $v_{j}^{-}=min(v_{ij})$, respectively.
Then the geometric distances $S_{j}^{+}$ and $S_{j}^{-}$ from $v_{ij}$ to the positive and negative ideal solutions can be computed, respectively, as
$S_{j}^{+}=\sqrt{\sum_{j=1}^{n}\left(v_{ij}-v_{j}^{+}\right)^{2}}$ (5)
and
$S_{j}^{-}=\sqrt{\sum_{j=1}^{n}\left(v_{ij}-v_{j}^{-}\right)^{2}}.$ (6)
The relative closeness $C_{i}^{*}$ for each agent was calculated as:
$C_{i}^{*}=\frac{S_{j}^{-}}{S_{j}^{+}+S_{j}^{-}}$ (7)
A higher $C_{i}^{*}$ signifies superior agent performance relative to peers on the assigned RP.
Acknowledgements
J.F.L. acknowledge supports from National Natural Science Foundation of China (Nos. 52394272, 22373022) and National Key Research and Development Program of China (No. 2023YFA0915300).
Supplementary information
We provide partial results for gravitational simulation mission along with examples of reports written by ASAs. (SI.pdf). We also provide partial AutoProgs and experimental result files for RP 1-3, main AI generating RP, and Multi-tier human RP. Some information has been obscured; readers can replace it with their own information and follow the instructions to run the AutoProgs (SI-data-1.zip and SI-data-2.zip). Additionally, we have created a demo video demonstrating the operation process of the automatic research system (SI-video.mp4).

### v2

@@
-Towards Fully Autonomous Research Powered by LLMs: Case Study on Simulations
-Abstract
-The advent of Large Language Models (LLMs) has created new opportunities for the automation of scientific research, spanning both experimental processes and computational simulations.
+Towards Fully Autonomous Research Powered by LLMs: Case Study on Simulations Zhihan Liu,a Yubo Chai,a Jianfeng Li,a
+ABSTRACT: The advent of Large Language Models (LLMs) has created new opportunities for the automation of scientific research, spanning both experimental processes and computational simulations.
 This study explores the feasibility of constructing an autonomous simulation agent (ASA) powered by LLM, through sophisticated API integration, to automate the entire research process, from experimental design, remote upload and simulation execution, data analysis, to report compilation.
@@
 Additionally, we discussed the intrinsic traits of ASAs in managing extensive tasks, focusing on self-validation mechanisms and the balance between local attention and global oversight.
+KEYWORDS: autonomous research, simulation, LLM, polymer physics
+footnotetext: a The State Key Laboratory of Molecular Engineering of Polymers, The Research Center of AI for Polymer Science Department of Macromolecular Science, Fudan University, Shanghai 200433, China; E-mail: $\ast$ lijf@fudan.edu.cn
 INTRODUCTION
-In recent years, the rapid development of artificial intelligence technology, particularly the emergence of large language models (LLMs), has brought revolutionary changes to scientific research[1, 2, 3, 4, 5].
-Since the launch of ChatGPT by OpenAI, LLMs such as GPT-4 and Anthropic’s Claude have demonstrated outstanding performance in natural language processing, text generation, and translation [6, 7].
-These AI models not only excel in traditional text processing tasks but also show significant potential in scientific research, gradually transforming conventional research methods and providing scientists with unprecedented tools and methodologies[8, 9, 10, 11].
+In recent years, the rapid development of artificial intelligence technology, particularly the emergence of large language models (LLMs), has brought revolutionary changes to scientific research1, 2, 3, 4, 5.
+Since the launch of ChatGPT by OpenAI, LLMs such as GPT-4 and Anthropic’s Claude have demonstrated outstanding performance in natural language processing, text generation, and translation 6, 7.
+These AI models not only excel in traditional text processing tasks but also show significant potential in scientific research, gradually transforming conventional research methods and providing scientists with unprecedented tools and methodologies8, 9, 10, 11.
 The application of artificial intelligence (AI) in the fields of chemistry and materials engineering is continually expanding.
-By integrating machine learning and automated experimental technologies, faster experimental design and data analysis can be achieved, leading to the discovery of new materials and the optimization of existing material properties in a shorter timeframe [12, 13, 14].
-In the domain of physical chemistry knowledge parsing, Meta’s Galactica assists researchers with tasks such as information organization, knowledge inference, and writing, showcasing the immense potential of large language models in scientific and professional domains [13].
-Meanwhile, DeepMind’s AlphaFold series of algorithms enable the prediction of structures for biological molecules, from proteins and nucleic acids to larger and more complex complexes [15, 16].
+By integrating machine learning and automated experimental technologies, faster experimental design and data analysis can be achieved, leading to the discovery of new materials and the optimization of existing material properties in a shorter timeframe 12, 13, 14.
+In the domain of physical chemistry knowledge parsing, Meta’s Galactica assists researchers with tasks such as information organization, knowledge inference, and writing, showcasing the immense potential of large language models in scientific and professional domains 13.
+Meanwhile, DeepMind’s AlphaFold series of algorithms enable the prediction of structures for biological molecules, from proteins and nucleic acids to larger and more complex complexes 15, 16.
 In the fields of chemical synthesis and drug design, AI aids in complex chemical synthesis planning, chemical property prediction, and the generation of new materials.
-Integrating LLMs with specialized chemical tools and databases can extend their capabilities beyond general language tasks [17, 18, 19, 20, 21].
+Integrating LLMs with specialized chemical tools and databases can extend their capabilities beyond general language tasks 17, 18, 19, 20, 21.
 Additionally, some studies have highlighted the crucial role of prompt engineering in guiding models to produce accurate and useful outputs.
-Techniques such as Chain of Thought, ReAct, and Tree of Thoughts help structure the model’s reasoning process, reduce errors, and improve the quality of generated results [20, 22, 23, 24].
+Techniques such as Chain of Thought, ReAct, and Tree of Thoughts help structure the model’s reasoning process, reduce errors, and improve the quality of generated results 20, 22, 23, 24.
 Recent research shows that continuous human-AI interaction can accelerate various stages of scientific research.
@@
 Some work has explored an AI-dominated research model, such as smart laboratories (SDLs) that integrate AI, laboratory automation, and robotics to conduct autonomous experiments.
-These systems can automatically adjust and optimize based on intelligent analysis and decision-making support from experimental results [25, 26, 27].
+These systems can automatically adjust and optimize based on intelligent analysis and decision-making support from experimental results 25, 26, 27.
 The ultimate goal for future AI scientists is to achieve full automation of scientific research processes, from experimental design to data analysis and conclusion generation.
@@
 Therefore, they exhibit different dynamic behaviors and scaling properties, with different scaling relationships between the mean square end-to-end distance and the number of chain segmentss N.
-This scaling relationship has been theoretically derived by Flory, De Gennes, and others [28, 29] and validated through simulations and experiment [30, 31].
+This scaling relationship has been theoretically derived by Flory, De Gennes, and others 28, 29 and validated through simulations and experiment 30, 31.
 This problem is fundamental in polymer physics and has been extensively studied.
@@
 In addition to the RPs designed for polymer conformation simulations, we have also crafted RPs for solar system planetary orbit simulations and asteroid trajectory simulations (Figure 1C and SI).
-Figure 1: A.
+Fig.
+1: A.
 Schematic diagram of the AI automation process for theoretical simulation research conducted in this study.
@@
 B.
-Diagram of the LLM dialogue.
+Diagram of the ASA dialogue.
 The human researcher only inputs the RP at the beginning, and then the ASA makes decisions based on the dialogue history, achieving a fully autonomous research mode.
@@
 Overall, this paper presents the first fully autonomous AI-conducted simulation research process powered by LLMs.
+As demonstrated in the demo video, simply submitting a research plan file via the command ’do AI4S_prj1.txt’ to ASA allows it to execute all subsequent tasks without any human intervention; ASA can even automatically upload programs to a remote server and perform a series of simulations—-something that cannot be achieved solely through interactions with a web-based LLM.
 By adopting this new research paradigm, we aim to explore the potential of AI in highly complex and data-intensive scientific research, breaking traditional human limitations and advancing the automation and intelligence of scientific research.
-Figure 2: Overview of RP 1-3.
-We designed three RPs, each containing multiple steps such as simulation, plotting, and report writing, and provided them to the ASA.
+Results
+Fig.
+2: Overview of RP 1-3.
+We designed three RPs related to polymer chain simulation, each containing multiple steps such as simulation, plotting, and report writing, and provided them to the ASA.
 RP 1 and RP 3 required the ASA to generate a complete Python program to simulate polymer chains, run the simulation locally, and produce a final report.
@@
 The provided simulation program (omitted in the figure) and remote server information were included in RP 2.
-Results
+Fig.
+3: Completion Rates of Various ASAs for RP 1-3.
+A.
+Table of LLM models used in this study.
+B.
+Statistics of the number of criteria met by each ASA over 20 experiments.
+We established seven criteria for RP 1-3 to measure mission completion rates and counted the number of times each ASA met these criteria across 20 experiments.
+Some zero results are not displayed.
+C.
+ASA Scoring.
+Relative scores for each ASA in RP 1-3 were calculated using the EWM and TOPSIS methods.
+D.
+ASA Generated Research Report Example.
+Demonstrates a Word report generated by ASA-GPT-4 for RP 1, including four sections, chain conformation diagrams and a scaling relation fit plot (also see SI).
 Automating Basic Simulations in Research
@@
 We provided RP 1 and 2 directly to the Master AI, adjusting the ASA to ensure each sub-task was handled by a distinct Subordinate AI without access to the Main AI’s or other Subordinate AIs’ conversation histories.
-The Main AI solely presents the AI RPs (Figure 4A) and received reports upon task completion.
+The Main AI solely presents the AI RPs (Figure S1A) and received reports upon task completion.
 We also employed GPT-4-turbo for this experiment.
@@
 RP 2 provides a random walk simulation program and details of a remote server, including hostname and username.
-Despite multiple attempts, the Main AI struggles to convey both the simulation program and server details accurately to the Subordinate AIs (Figure 4B).
+Despite multiple attempts, the Main AI struggles to convey both the simulation program and server details accurately to the Subordinate AIs (Figure S1B).
 We attempted to instruct the Main AI to include all relevant information in the AI RPs, but this did not notably enhance the accuracy of information transmission.
@@
 Each execution generates an Agent AI to fulfill RP 1 contained in p1.txt, while the Primary AI collects all generated files and conducts result analysis.
-The detailed content of RP 4 is depicted in Figure 4C.
+The detailed content of RP 4 is depicted in Figure S1C.
 Using ASA-GPT-4-Turbo, we conducted tests on RP 4.
@@
 This illustrates the remarkable capability of multi-tier RP structures in handling large-scale missions.
-Figure 3: Completion Rates of Various ASAs for RP 1-3.
-A.
-Table of LLM models used in this study.
-B.
-Statistics of the number of criteria met by each ASA over 20 experiments.
-We established seven criteria for RP 1-3 to measure mission completion rates and counted the number of times each ASA met these criteria across 20 experiments.
-Some zero results are not displayed.
-C.
-ASA Scoring.
-Relative scores for each ASA in RP 1-3 were calculated using the EWM and TOPSIS methods.
-D.
-ASA Generated Research Report Example.
-Demonstrates a Word report generated by ASA-GPT-4 for RP 1, including four sections, chain conformation diagrams and a scaling relation fit plot (also see SI).
-Figure 4: A.
-Example of Main AI’s Response.
-The Main AI first breaks down the overall mission into sub-tasks, then generates AI RPs beginning with ’< < < prompt’ and ending with ’end > > >’.
-The AI-generated RPs are then passed to subordinate AIs for execution.
-B.
-Example of Faulty and Incomplete AI RP.
-The black box highlights where the AI RP fails to provide necessary information such as username and password for connecting to the remote server, and the red box highlights where the AI RP includes irrelevant information.
-C.
-Overview of RP 4.
-The command ’python AI4SCI _VE13.py -s p1.txt -n i’ generates an Agent AI to complete the mission of RP 1, where AI4SCI _VE13.py is the API automation execution program, p1.txt is the text file containing the RP, and i denotes the $i^{th}$ execution of the command.
 Discussion
-The experiments demonstrate that LLMs, with their robust coding ability [24, 32, 33, 34], can be directly employed in developing agents to autonomously complete complex research missions when aided by proper API programming.
+The experiments demonstrate that LLMs, with their robust coding ability 24, 32, 33, 34, can be directly employed in developing agents to autonomously complete complex research missions when aided by proper API programming.
 ASA-GPT-4o and ASA-GPT-4-turbo, notably, exhibits near-perfect completion rates for simpler missions, highlighting their potential.
 Multi-tier RP designs effectively handle more complex missions, though accurate information transmission remains challenging for such RPs.
+We emphasize that these simulation tasks, particularly those involving remote server execution, cannot be accomplished through interactions with web-based LLMs alone.
 Enhancements in RP design and AI coordination are essential for fully harnessing AI’s potential in automating scientific research.
-Below, we discuss several problems existing in the current reaserch.
+Below, we discuss the insights and existing problems identified in the current research.
+Recommended Practices for Fully Autonomous Research
+ASA achieves fully autonomous research through continuous interaction with the LLM according to a specific logical design guided by the research plan (RP) provided by the researchers.
+We have identified several key practices in ASA logic design and RP design to enhance execution success rates.
+  • Structure agent responses for efficient code and task extraction.
+Require the LLM to produce structured outputs, enabling ASA to accurately extract Python code, subtasks, and progress indicators.
+This improves code execution success rates and ensures smooth mission progression.
+  • Incorporate agent-led code review and debugging.
+While the LLM is powerful in code generation, it may struggle with complex programs on the first attempt.
+By integrating self-review and debugging process, the LLM can generate correct and task-compliant code within several dialogue rounds.
+  • Ensure key outputs at each subtask step for continuity.
+Long missions often need to be divided into subtasks.
+To maintain coherence, prompt the LLM to consider previous outputs, generate key results for each step and input key results into the proceeding round of dialogue, facilitating the continuation of subsequent tasks and allowing the agent to monitor overall progress.
+  • Provide detailed information for complex operations like remote server connections.
+Providing sufficient information in RP can improve mission success rates.
+In addition to necessary numeric/textual information, sometimes offering specific methods can guide the LLM to write correct programs more efficiently.
+In summary, while the LLM has strong generative capabilities, its unpredictable outputs can complicate automation.
+ASA, based on Python’s AutoProg, though less intelligent, is reliable when logic is well-designed.
+By effectively structuring logic, ASA can oversee the LLM’s output, enhancing process reliability and achieving fully autonomous research.
+Fig.
+4: Recommended practices for fully autonomous research.
+By conducting iterative dialogues with the LLM adhering to a specific design logic, the success rate and efficiency of the automated process are enhanced.
+The diagrams on the left and right illustrate scenarios without and with the implementation of these strategies, respectively.
 Common Issues in Sequential Task
@@
 In another example, for simulating self-avoiding walks, ASAs often incorrectly extend the random walk strategy by merely checking distances between points, resulting in high-energy state samples with large statistical errors.
-Correct sampling should use importance sampling to generate statistically significant conformations[35].
+Correct sampling should use importance sampling to generate statistically significant conformations35.
 To avoid these issues in research, reliable solutions should be provided in the RPs, or relevant domain knowledge should be used for fine-tuning the LLMs to enhance their understanding.
@@
 This phenomenon may be associated with the LLM’s propensity to depend heavily on past dialogue content.
-Research has documented instances where LLMs utilize context to deliver responses preferred by users, indicating an excessive reliance on previous information [36, 37, 38].
+Research has documented instances where LLMs utilize context to deliver responses preferred by users, indicating an excessive reliance on previous information 36, 37, 38.
 While this capability allows LLMs to adapt responses based on short dialogue histories, it can also lead to persistent errors.
@@
 This study demonstrates the potential of LLM powered research agent ASA to autonomously conduct scientific research missions, using a polymer theory simulation problem and a more challenging asteroid-trajectory prediction problem as model problems.
-This process, encompassing simulation coding, remote execution, data analysis, and report generation, represents the first fully autonomous AI-directed simulation research.
+This process, encompassing simulation coding, remote execution, data analysis, and report generation, represents the first fully autonomous AI-directed simulation research: once the research plan is made, only a single command (e.g., ‘do prj.txt’) is required to initiate and complete the process.
 The notable success herein can be largely attributed to the sophisticated API programming.
@@
 These capabilities underscore the readiness of ASAs to independently manage long-term scientific research missions, offering a new perspective on AI applications in science as well as in chemical engineering.
-Method
+Methods
 API automation Program (AutoProg)
@@
 For quantitative analysis and scoring, we deployed the Entropy Weight Method (EWM) and TOPSIS (Technique for Order of Preference by Similarity to Ideal Solution).
-The EWM, a well-established approach for determining the relative significance of criteria in decision-making [39], was utilized to assign weights based on data variability.
+The EWM, a well-established approach for determining the relative significance of criteria in decision-making 39, was utilized to assign weights based on data variability.
 This ensures criteria with higher variability, and thus more informational value, receive higher weights in the final evaluation.
-TOPSIS, originally formulated by Hwang and Yoon [40], was subsequently applied.
+TOPSIS, originally formulated by Hwang and Yoon 40, was subsequently applied.
 This method computes the geometric distance between each agent’s performance and both ideal and anti-ideal solutions, facilitating a comparative assessment of ASA efficacy.
@@
 A higher $C_{i}^{*}$ signifies superior agent performance relative to peers on the assigned RP.
+ASSOCIATED CONTENT
+Supporting Information
+The following Supporting Information is available free of charge.
+Partial results for gravitational simulation mission along with examples of reports written by ASAs (SI.pdf); partial AutoProgs and experimental result files for RP 1-3, main AI generating RP, and Multi-tier human RP (Some information has been obscured; readers can replace it with their own information and follow the instructions to run the AutoProgs) (SI-data-1.zip and SI-data-2.zip); a demo video demonstrating the operation process of the automatic research system (SI-video.mp4).
+AUTHOR INFORMATION
+Corresponding Author
+Jianfeng Li - The State Key Laboratory of Molecular Engineering of Polymers, The Research Center of AI for Polymer Science
+Department of Macromolecular Science, Fudan University, Shanghai 200433, China; Email: lijf@fudan.edu.cn
+Authors
+Zhihan Liu - The State Key Laboratory of Molecular Engineering of Polymers, The Research Center of AI for Polymer Science
+Department of Macromolecular Science, Fudan University, Shanghai 200433, China
+Yubo Chai - The State Key Laboratory of Molecular Engineering of Polymers, The Research Center of AI for Polymer Science
+Department of Macromolecular Science, Fudan University, Shanghai 200433, China
+Author Contributions
+Z.H.L. was the primary author of the paper. Y.B.C. was responsible for part of the data analysis and statistics. J.F.L. originally initiated and supervised the project.
+Notes
+The authors declare no competing financial interest.
 ACKNOWLEDGEMENTS
 J.F.L. acknowledge supports from National Natural Science Foundation of China (Nos. 52394272, 22373022) and National Key Research and Development Program of China (No. 2023YFA0915300).
-Supplementary information
-We provide partial results for gravitational simulation mission along with examples of reports written by ASAs.
-(SI.pdf).
-We also provide partial AutoProgs and experimental result files for RP 1-3, main AI generating RP, and Multi-tier human RP.
-Some information has been obscured; readers can replace it with their own information and follow the instructions to run the AutoProgs (SI-data-1.zip and SI-data-2.zip).
-Additionally, we have created a demo video demonstrating the operation process of the automatic research system (SI-video.mp4).
+Notes and references
+  • Devlin et al.
+2019 Devlin, J.; Chang, M.-W.; Lee, K.; Toutanova, K.
+BERT: Pre-training of Deep Bidirectional Transformers for Language Understanding.
+Proceedings of the 2019 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long and Short Papers).
+Minneapolis, Minnesota, 2019; pp 4171–4186.
+  • Raffel et al.
+2020 Raffel, C.; Shazeer, N.; Roberts, A.; Lee, K.; Narang, S.; Matena, M.; Zhou, Y.; Li, W.; Liu, P.
+J.
+Exploring the Limits of Transfer Learning with a Unified Text-to-Text Transformer.
+Journal of Machine Learning Research 2020, 21, 1–67.
+  • Floridi and Chiriatti 2020 Floridi, L.; Chiriatti, M. GPT-3: Its Nature, Scope, Limits, and Consequences. Minds and Machines 2020, 30, 681–694.
+  • Chowdhery et al. 2022 Chowdhery, A. et al. PaLM: Scaling Language Modeling with Pathways. Journal of Machine Learning Research 2022, 24, 240:1–240:113.
+  • Thoppilan et al. 2022 Thoppilan, R. et al. LaMDA: Language Models for Dialog Applications. arXiv:2201.08239, 2022.
+  • OpenAI 2023 OpenAI GPT-4 Technical Report. https://cdn.openai.com/papers/gpt-4.pdf, 2023; Retrieved June 24, 2024.
+  • Anthropic 2023 Anthropic Model Card and Evaluations for Claude Models.
+https://www-cdn.anthropic.com/files/4zrzovbb/website/bd2a28d2535bfb0494cc8e2a3bf135d2e7523226.pdf, 2023; Retrieved June 24, 2024.
+  • Ramesh et al.
+2021 Ramesh, A.; Pavlov, M.; Goh, G.; Gray, S.; Voss, C.; Radford, A.; Chen, M.; Sutskever, I.
+Zero-shot text-to-image generation.
+International conference on machine learning.
+2021; pp 8821–8831.
+  • Bommasani et al. 2022 Bommasani, R. et al. On the Opportunities and Risks of Foundation Models. arXiv:2108.07258, 2022.
+  • Chang et al. 2024 Chang, Y. et al. A Survey on Evaluation of Large Language Models. ACM Transactions on Intelligent Systems and Technology 2024, 15.
+  • Kasneci et al. 2023 Kasneci, E. et al. ChatGPT for good? On opportunities and challenges of large language models for education. Learning and Individual Differences 2023, 103, 102274.
+  • Himanen et al. 2019 Himanen, L.; Geurts, A.; Foster, A. S.; Rinke, P. Data-Driven Materials Science: Status, Challenges, and Perspectives. Advanced Science 2019, 6, 1900808.
+  • Taylor et al. 2022 Taylor, R.; Kardas, M.; Cucurull, G.; et al. Galactica: A Large Language Model for Science. arXiv:2211.09085, 2022.
+  • Wang et al. 2023 Wang, H.; Fu, T.; Du, Y.; et al. Scientific discovery in the age of artificial intelligence. Nature 2023, 620, 47–60.
+  • Evans et al. 2021 Evans, R.; O’Neill, M.; Pritzel, A.; et al. Protein complex prediction with AlphaFold-Multimer. bioRxiv:2021.10.04.463034, 2021.
+  • Abramson et al. 2024 Abramson, J.; Adler, J.; Dunger, J.; et al. Accurate structure prediction of biomolecular interactions with AlphaFold 3. Nature 2024, 630, 493–500.
+  • Brown et al. 2019 Brown, N.; Fiscato, M.; Segler, M.; et al. GuacaMol: Benchmarking Models for de Novo Molecular Design. Journal of Chemical Information and Modeling 2019, 59, 1096–1108.
+  • Stokes et al. 2020 Stokes, J. M.; Yang, K.; Swanson, K.; et al. A deep learning approach to antibiotic discovery. Cell 2020, 180, 688–702.
+  • Boiko et al. 2023 Boiko, D. A.; MacKnight, R.; Kline, B.; et al. Autonomous chemical research with large language models. Nature 2023, 624, 570–578.
+  • Bran et al. 2024 Bran, M.; Cox, A.; Schilter, S.; et al. Augmenting large language models with chemistry tools. Nature Machine Intelligence 2024, 6, 525–535.
+  • Kang and Kim 2024 Kang, Y.; Kim, J.
+ChatMOF: an artificial intelligence system for predicting and generating metal-organic frameworks using large language models.
+Nature Communications 2024, 15, 4705.
+  • Jablonka et al. 2024 Jablonka, K. M.; Schwaller, P.; Ortega-Guerrero, A.; et al. Leveraging large language models for predictive chemistry. Nature Machine Intelligence 2024, 6, 161–169.
+  • Schulhoff et al. 2024 Schulhoff, S.; Ilie, M.; Balepur, N.; et al. The Prompt Report: A Systematic Survey of Prompting Techniques. arXiv:2406.06608, 2024.
+  • Zhou et al. 2024 Zhou, Z.; Ning, X.; Hong, K.; et al. A Survey on Efficient Inference for Large Language Models. arXiv:2404.14294, 2024.
+  • Abolhasani and Kumacheva 2023 Abolhasani, M.; Kumacheva, E. The rise of self-driving labs in chemical and materials sciences. Nature Synthesis 2023, 2, 483–492.
+  • Szymanski et al. 2023 Szymanski, N. J.; Rendy, B.; Fei, Y.; et al. An autonomous laboratory for the accelerated synthesis of novel materials. Nature 2023, 624, 86–91.
+  • Volk et al.
+2023 Volk, A.
+A.; Epps, R.
+W.; Yonemoto, D.
+T.; et al.
+AlphaFlow: autonomous discovery and optimization of multi-step chemistry using a self-driven fluidic lab guided by reinforcement learning.
+Nature Communications 2023, 14, 1403.
+  • Flory 1969 Flory, P. J. Statistical Mechanics of Chain Molecules; Interscience Publishers, 1969.
+  • De Gennes 1979 De Gennes, P. G. Scaling concepts in polymer physics; Cornell University Press, 1979.
+  • Madras and Slade 2013 Madras, N.; Slade, G. The self-avoiding walk; Birkhäuser Boston, 2013.
+  • Burchard 1983 Burchard, W. Static and Dynamic Light Scattering from Branched Polymers and Biopolymers. Advances in Polymer Science 1983, 48, 1–124.
+  • Fang et al. 2023 Fang, C.; Miao, N.; Srivastav, S.; et al. Large Language Models for Code Analysis: Do LLMs Really Do Their Job? arXiv:2310.12357, 2023.
+  • Zhong and Wang 2023 Zhong, L.; Wang, Z. Can ChatGPT replace StackOverflow? A Study on Robustness and Reliability of Large Language Model Code Generation. arXiv:2308.10335, 2023.
+  • Murr et al. 2023 Murr, L.; Grainger, M.; Gao, D. Testing LLMs on Code Generation with Varying Levels of Prompt Specificity. arXiv:2311.07599, 2023.
+  • Madras and Sokal 1988 Madras, N.; Sokal, A. D. The pivot algorithm: A highly efficient Monte Carlo method for the self-avoiding walk. Journal of Statistical Physics 1988, 50, 109–186.
+  • Brown et al. 2020 Brown, T. B.; Mann, B.; Ryder, N. Language Models are Few-Shot Learners. arXiv:2005.14165, 2020.
+  • Gao et al. 2020 Gao, T.; Fisch, A.; Chen, D. Making Pre-trained Language Models Better Few-shot Learners. arXiv:2012.15723, 2020.
+  • Wei et al.
+2022 Wei, J.; Bosma, M.; Zhao, V.
+Y.; Guu, K.; Yu, A.
+W.; Lester, B.; Du, N.; Dai, A.
+M.; Le, Q.
+V.
+Finetuned Language Models are Zero-Shot Learners.
+The Tenth International Conference on Learning Representations, ICLR 2022, Virtual Event, April 25-29, 2022.
+2022.
+  • Ravindran et al. 1987 Ravindran, A.; Phillips, D. T.; Solberg, J. J. Operations Research Principles & Practice, 2nd ed.; John Wiley: New York, 1987.
+  • Hwang and Yoon 1981 Hwang, C. L.; Yoon, K. Multiple Attribute Decision Making; Lecture Notes in Economics and Mathematical Systems; Springer: Berlin, Heidelberg, 1981; Vol. 186.

### v3

@@
-Towards Fully Autonomous Research Powered by LLMs: Case Study on Simulations Zhihan Liu,a Yubo Chai,a Jianfeng Li,a
-ABSTRACT: The advent of Large Language Models (LLMs) has created new opportunities for the automation of scientific research, spanning both experimental processes and computational simulations.
-This study explores the feasibility of constructing an autonomous simulation agent (ASA) powered by LLM, through sophisticated API integration, to automate the entire research process, from experimental design, remote upload and simulation execution, data analysis, to report compilation.
-Using a simulation problem of polymer chain conformations as a case study, we assessed the performance of ASAs powered by different LLMs including GPT-4-Turbo.
-Our findings revealed that ASA-GPT-4o achieved near-flawless execution on designated research missions, underscoring the potential of LLMs to manage complete scientific investigations autonomously.
-The outlined automation can be iteratively performed up to twenty cycles without human intervention, illustrating the potential of LLMs for large-scale autonomous research endeavors.
-Additionally, we discussed the intrinsic traits of ASAs in managing extensive tasks, focusing on self-validation mechanisms and the balance between local attention and global oversight.
+Toward Automated Simulation Research Workflow through LLM Prompt Engineering Design Zhihan Liu,a Yubo Chai,a Jianfeng Li,a
+ABSTRACT: The advent of Large Language Models (LLMs) has created new opportunities for the automation of scientific research spanning both experimental processes and computational simulations.
+This study explores the feasibility of constructing an autonomous simulation agent (ASA) powered by LLMs through prompt engineering and automated program design to automate the entire simulation research process according to a human-provided research plan.
+This process includes experimental design, remote upload and simulation execution, data analysis, and report compilation.
+Using a well-studied simulation problem of polymer chain conformations as a test case, we assessed the long-task completion and reliability of ASAs powered by different LLMs, including GPT-4o, Claude-3.5, etc.
+Our findings revealed that ASA-GPT-4o achieved near-flawless execution on designated research missions, underscoring the potential of methods like ASA to achieve automation in simulation research processes to enhance research efficiency.
+The outlined automation can be iteratively performed for up to 20 cycles without human intervention, illustrating the potential of ASA for long-task workflow automation.
+Additionally, we discussed the intrinsic traits of ASA in managing extensive tasks, focusing on self-validation mechanisms, and the balance between local attention and global oversight.
 KEYWORDS: autonomous research, simulation, LLM, polymer physics
@@
 INTRODUCTION
-In recent years, the rapid development of artificial intelligence technology, particularly the emergence of large language models (LLMs), has brought revolutionary changes to scientific research1, 2, 3, 4, 5.
-Since the launch of ChatGPT by OpenAI, LLMs such as GPT-4 and Anthropic’s Claude have demonstrated outstanding performance in natural language processing, text generation, and translation 6, 7.
-These AI models not only excel in traditional text processing tasks but also show significant potential in scientific research, gradually transforming conventional research methods and providing scientists with unprecedented tools and methodologies8, 9, 10, 11.
-The application of artificial intelligence (AI) in the fields of chemistry and materials engineering is continually expanding.
-By integrating machine learning and automated experimental technologies, faster experimental design and data analysis can be achieved, leading to the discovery of new materials and the optimization of existing material properties in a shorter timeframe 12, 13, 14.
-In the domain of physical chemistry knowledge parsing, Meta’s Galactica assists researchers with tasks such as information organization, knowledge inference, and writing, showcasing the immense potential of large language models in scientific and professional domains 13.
-Meanwhile, DeepMind’s AlphaFold series of algorithms enable the prediction of structures for biological molecules, from proteins and nucleic acids to larger and more complex complexes 15, 16.
-In the fields of chemical synthesis and drug design, AI aids in complex chemical synthesis planning, chemical property prediction, and the generation of new materials.
-Integrating LLMs with specialized chemical tools and databases can extend their capabilities beyond general language tasks 17, 18, 19, 20, 21.
-Additionally, some studies have highlighted the crucial role of prompt engineering in guiding models to produce accurate and useful outputs.
-Techniques such as Chain of Thought, ReAct, and Tree of Thoughts help structure the model’s reasoning process, reduce errors, and improve the quality of generated results 20, 22, 23, 24.
-Recent research shows that continuous human-AI interaction can accelerate various stages of scientific research.
-However, this human-led, AI-assisted model still limits the speed and depth of scientific research.
-Some work has explored an AI-dominated research model, such as smart laboratories (SDLs) that integrate AI, laboratory automation, and robotics to conduct autonomous experiments.
-These systems can automatically adjust and optimize based on intelligent analysis and decision-making support from experimental results 25, 26, 27.
-The ultimate goal for future AI scientists is to achieve full automation of scientific research processes, from experimental design to data analysis and conclusion generation.
-This end-to-end automation is set to become a major development in chemical research and a primary objective in chemical engineering, including both experimental and simulation aspects.
-This paper explores the AI automation paradigm in polymer theory simulation research.
-In traditional simulation research, a supervisor typically discusses and plans the research project with students.
-The students then write and modify the simulation program based on physical models, run simulations under different parameter conditions on remote servers, collect data, analyze results, and write reports.
-This paper envisions a new research paradigm where, after the research has been planned, the subsequent tasks are solely carried out by AI.
-Using an autonomous simulation agent (ASA) powered by LLM, after human researchers input a research plan (RP), the ASA will then take on the role of the student to complete the entire research process.
-Humans only need to review the AI’s results at the end, thus achieving a research mode without human intervention.
-To test this research paradigm, we systematically attempted ASAs powered by different LLMs on a moderately complex simulation problem in polymer physics.
-This involved an AI fully autonomously coding the simulation program, remote uploading and executing simulations, data organization, plotting, and scientific report writing.
-Random walk and self-avoiding walk models are of significant physical importance in describing the spatial configuration of polymer chains.
-The random walk model assumes that each ‘step’ of the chain is independent and random, while the self-avoiding walk model considers the volume exclusion effects between chain segments.
-Therefore, they exhibit different dynamic behaviors and scaling properties, with different scaling relationships between the mean square end-to-end distance and the number of chain segmentss N.
-This scaling relationship has been theoretically derived by Flory, De Gennes, and others 28, 29 and validated through simulations and experiment 30, 31.
-This problem is fundamental in polymer physics and has been extensively studied.
-However, for AI, it still presents a novel challenge.
-The difficulty of predicting the scaling exponent makes it neither too easy nor prohibitively difficult, thus serving as an excellent model problem for testing AI capabilities.
-We prompted the AI to write simulation programs for both random walk and self-avoiding walk models, compute the mean square end-to-end distance for different chain segment numbers N, fit the results to $\left\langle R^{2}\right\rangle\propto N^{\nu}$ , determine the scaling exponent $\nu$, plot the data, and write a complete scientific report.
-The ASA leverages an API automation program (AutoProg) to enable multi-round text interactions with LLMs via API (Figure 1).
-Initially, the AutoProg provides an RP witten by human researchers to an LLM, and the LLM will break the plan down into multiple sub-tasks and write Python code to complete the first sub-task in its response.
-The AutoProg then extracts the Python code, executes it, and returns the execution results and any errors along with the conversation history back to the LLM.
-The LLM then decide whether to modify the Python code or to proceed to the following sub-tasks.
-This process repeats until the LLM determines that the mission is complete, at which point the AutoProg stops.
-Upon mission completion, humans evaluate its success based on criteria such as correct simulation, plotting, etc.
-The ASA we developed is a general system (minor adjustments may be required in the AutoProg for different LLM APIs) that requires only a change in the original RP to be applicable to any simulation problem involving Python programming.
-In addition to the RPs designed for polymer conformation simulations, we have also crafted RPs for solar system planetary orbit simulations and asteroid trajectory simulations (Figure 1C and SI).
+In recent years, the application of artificial intelligence (AI) in scientific research has gained significant traction.
+AI technologies, through data analysis, pattern recognition, and predictive capabilities, offer researchers unprecedented support, enhancing both efficiency and accuracy.
+AI is utilized in chemical synthesis, drug design, and materials science, accelerating the achievement of valuable research outcomes.
+Meta’s Galactica supports researchers in organizing information, inferring knowledge, and writing, highlighting the potential of AIs in scientific contexts.1 DeepMind’s AlphaFold algorithms enable the prediction of structures for biological molecules, from proteins and nucleic acids to larger and more complex complexes.2, 3 Machine Learning methods not only enhance the understanding and predictive capabilities of complex systems but also offer new approaches to tackling problems that are difficult for traditional computational methods to handle.4, 5, 6, 7, 8, 9 Integrating machine learning with automated experimental techniques enables faster experimental design and data analysis, leading to quicker discoveries of new materials and improvements in existing material properties.10, 11
+Among various AI technologies, tools based on large language models (LLMs) such as GPT and Claude have demonstrated superior capabilities in generating and understanding complex texts across a range of natural language processing tasks.12, 13 In addition, LLMs can also generate corresponding code based on text descriptions, significantly simplifying programming tasks.14, 15 Research shows that LLMs excel in handling complex logic, code completion, and debugging, making them valuable assistants in software development and data analysis.16, 17, 18, 19, 20 In the field of scientific research, the application prospects of LLMs are increasingly evident, gradually transforming traditional research methods and providing scientists with unprecedented tools and methodologies.21, 22, 23, 24 For example, integrating LLMs with specialized chemical tools and databases can extend their capabilities beyond general language tasks.25, 26, 27, 28, 29 Researchers have also utilized LLMs to automate literature reviews30, 31, 32, 33 and data analysis34, 35, 36 and assist in experimental design and species screening 37, 38, 39, 29, 40 to improve research efficiency.
+Additionally, some studies highlighted the crucial role of prompt engineering in guiding models to produce accurate and useful outputs.
+Techniques such as Chain of Thought, ReAct, and Tree of Thoughts help structure the model’s reasoning process, reduce errors, and improve the quality of generated results.28, 41, 42, 43, 44, 45, 46 These advancements allow researchers to dedicate more time to innovative thinking and theoretical exploration rather than repetitive and inefficient experimental operations.
+The aforementioned studies show that continuous human-AI interaction can accelerate various stages of scientific research.
+However, the current model of these studies mainly involves interaction between humans and AI.
+By utilizing the robust capabilities of large language models (LLMs), we aspire to gradually reduce human involvement in the research process, aiming to enable AI to independently complete end-to-end workflows, thus further enhancing research efficiency.
+Taking simulation research as an example, in traditional simulation research processes, researchers develop a research plan (RP) that specifies the simulation methods to be used, how to write programs or use simulation software, how to build models, and which parameters to test.
+Following the establishment of the RP, a significant amount of experimentation is carried out to gather data, which are then analyzed to draw scientific conclusions.
+After the RP is determined, the subsequent steps are largely standardized and present opportunities for automation.
+Therefore, we intend to test whether, by providing a detailed textual RP, the advanced capabilities of LLMs can allow AI to successfully and automatically complete an entire simulation research process without human intervention.
+This approach could streamline the research workflow and lead to more efficient outcomes.
+With their outstanding language understanding and code generation capabilities, LLMs can produce rich and varied content based on prompts.
+However, they cannot guarantee the accuracy of the generated code in a single attempt, and due to the length and complexity of tasks described in the RP, LLMs may overlook subtasks.
+Additionally, as language models, LLMs do not have the ability to control local computers to execute programs.
+To address these challenges and achieve an automated workflow, we developed an automated system that can identify and run the code generated by the LLM on a local machine.
+We also used prompt engineering to guide the LLM in producing structured and accurate responses.
+Through multiple rounds of interaction, we were able to instruct the LLM to debug program errors, check the completeness of task execution, and make the necessary adjustments.
+This approach resulted in the creation of an end-to-end Autonomous Simulation Agent (ASA).
+As demonstrated in the demo video, simply submitting an RP txt file via the command “do AI4PS_PRJ1.txt” to ASA allows it to execute all subsequent tasks without any human intervention; ASA can even automatically upload programs to a remote server and perform a series of simulations—something that cannot be achieved solely through interactions with a web-based LLM.
+We chose a problem in polymer physics with known outcomes to evaluate the feasibility and reliability of ASA in automating research workflows.
+Detailed RPs were provided to the ASA, which involved the AI fully autonomously coding simulation programs, remotely uploading and executing simulations, organizing data, plotting results, and writing scientific reports.
+We also designed several evaluation criteria to determine if the ASA could correctly and completely perform each step according to RP without human intervention.
 Fig.
-1: A.
-Schematic diagram of the AI automation process for theoretical simulation research conducted in this study.
-The human researcher writes the research plan (RP) for the theoretical simulation and provides it to the autonomous simulation agent (ASA), which is developed by sophisticated API programming (see Method section).
-The ASA first generates the simulation program, then uploads it to a remote server, performs simulation calculations with different parameter conditions, collects data, and finally writes a report based on the simulation results.
-B.
-Diagram of the ASA dialogue.
-The human researcher only inputs the RP at the beginning, and then the ASA makes decisions based on the dialogue history, achieving a fully autonomous research mode.
-C.
-Two typical simulation problems studied in this paper: the chain conformation simulation of a random walk and the gravitational model simulation of asteroid orbit prediction (See SI).
-We compared the performance of ASAs powered by various LLMs, including GPT-4o, and explored common characteristics and issues when LLMs tackle long tasks.
-These findings provide insights for researchers to better improve and utilize LLMs for AI-driven scientific research.
-Overall, this paper presents the first fully autonomous AI-conducted simulation research process powered by LLMs.
-As demonstrated in the demo video, simply submitting a research plan file via the command ’do AI4S_prj1.txt’ to ASA allows it to execute all subsequent tasks without any human intervention; ASA can even automatically upload programs to a remote server and perform a series of simulations—-something that cannot be achieved solely through interactions with a web-based LLM.
-By adopting this new research paradigm, we aim to explore the potential of AI in highly complex and data-intensive scientific research, breaking traditional human limitations and advancing the automation and intelligence of scientific research.
+1: (A) Schematic diagram of the AI automation process for theoretical simulation research conducted in this study.
+The human researcher writes the research plan (RP) for the theoretical simulation and provides it to the autonomous simulation agent (ASA), which is developed by prompt engineering and automated program design (see the Methods section).
+The ASA first generates the simulation program and then uploads it to a remote server, performs simulation calculations with different parameter conditions, collects data, and finally writes a report based on the simulation results.
+(B) Diagram of the ASA dialogue.
+The human researcher only inputs the RP at the beginning, and then, the ASA makes actions based on the given RP and dialogue history, achieving an automated simulation research process.
+(C) Two simulation problems are used to test the ASA method in this paper: the chain conformation simulation of a random walk and the gravitational model simulation of asteroid orbit prediction (see the SI).
+Furthermore, we compared the performance of ASAs powered by various LLMs, including GPT-4o, GPT-4-Turbo, Claude-3.5, and so forth, and explored common characteristics and issues when ASAs tackle long tasks.
+These findings provide insights for researchers to improve and utilize LLMs better for AI-driven scientific research.
+Overall, this study represents the first exploration of an AI-driven, end-to-end simulation research process guided by an RP provided by humans.
+By using known problems and predefined evaluation criteria, we tested whether methods like ASA, constructed through an LLM prompt engineering and automated program, can accurately and completely execute the entire simulation research process by simply reading the textual information in RP.
+Through this, we aim to break traditional human limitations and advance the automation and efficiency of scientific research, particularly in data-intensive simulation studies (Figure 1).
 Results
@@
 3: Completion Rates of Various ASAs for RP 1-3.
-A.
-Table of LLM models used in this study.
-B.
-Statistics of the number of criteria met by each ASA over 20 experiments.
-We established seven criteria for RP 1-3 to measure mission completion rates and counted the number of times each ASA met these criteria across 20 experiments.
+(A) Table of LLM models tested in this study.
+(B) Statistics of the number of criteria met by each ASA over 20 experiments.
+We established seven criteria for each RP to measure mission completion rates and counted the number of times each ASA met these criteria across 20 experiments.
 Some zero results are not displayed.
-C.
-ASA Scoring.
-Relative scores for each ASA in RP 1-3 were calculated using the EWM and TOPSIS methods.
-D.
-ASA Generated Research Report Example.
-Demonstrates a Word report generated by ASA-GPT-4 for RP 1, including four sections, chain conformation diagrams and a scaling relation fit plot (also see SI).
+(C) ASA Scoring.
+Relative scores for each ASA in RP 1-3 were calculated using the EWM and TOPSIS methods (see the Methods section).
+(D) ASA Generated Research Report Example.
+Demonstrates screen cuts for a Word report generated by ASA-GPT-4o for RP 1, including four sections, chain conformation diagrams and a scaling relation fit plot ( see the full text in the SI).
+Automating Test Simulation Problems
+We tested the reliability and stability of the ASA approach using a well-studied polymer chain conformation problem.
+Random walking and self-avoiding walking models are of significant physical importance in describing the spatial configuration of polymer chains.
+The random walk model assumes that each “step” of the chain is independent and random, while the self-avoiding walk model considers the volume exclusion effects between chain segments.
+Therefore, they exhibit different dynamic behaviors and scaling properties, with different scaling relationships between the mean square end-to-end distance and the number of chain segments N.
+This scaling relationship has been theoretically derived by Flory, De Gennes, and others 47, 48 and validated through simulations and experiments.49, 50 This problem is fundamental to polymer physics and has been extensively studied.
+However, AI still presents a novel challenge.
+The difficulty of predicting the scaling exponent makes it neither too easy nor prohibitively difficult, thus serving as an excellent model problem for testing AI capabilities.
+We hope that the AI can write simulation programs for both random walk and self-avoiding walk models, compute the mean square end-to-end distance $\left\langle R^{2}\right\rangle$ for different chain segment numbers N, fit the results to $\left\langle R^{2}\right\rangle\propto N^{\nu}$, determine the scaling exponent $\nu$, plot the data, and write a complete scientific report.
+We designed three simulation missions related to polymer chain conformation modeling to demonstrate and test the ability of ASAs powered by different LLMs to autonomously complete the entire research process based on the given research plans (RPs).
+Each mission includes multiple subtasks, covering the full spectrum, from conducting experiments to analyzing data and writing reports.
+The RPs for these missions are listed in Figure 2.
+We established several evaluation criteria based on the requirements of RP 1-3, such as whether graphs and plots were generated and whether the derived scaling relationships were correct, to assess task completion.
+We tested nine ASAs powered by different LLMs, including GPT-4o, conducting 20 trials for RP 1-3 and recording the achievement of the evaluation criteria across these trials, as shown in Figure 3.
+Subsequent sections will analyze the performance of ASAs on RPs 1-3.
+Despite well-studied chain conformation simulations being able to effectively demonstrate the reliability of ASA in automatically executing simulation research procedures based on RPs, to demonstrate that ASA can solve problems across different scientific domains, we also designed RPs for more challenging gravitational simulation problems corresponding to RP S1 and RP S2 (Figure S2 and 1C).
+The mission of RP S2 is to evaluate the risk posed by a meteorite (or asteroid) to Earth, given its initial position and velocity and write a report.
+This mission is challenging even for graduates of physics departments.
+By simply inputting RP S2 into the same ASA used for chain conformation problems, the ASA will automatically acquire the necessary parameters of the solar system, code the simulation program to calculate the asteroid’s trajectory, and ultimately write the report.
+This result underscores the effectiveness of our prompt engineering and automated program design strategy (see the Methods section) and strongly demonstrates that ASA can address problems in entirely different fields.
 Automating Basic Simulations in Research
-We designed three simulation missions related to polymer chain conformation modeling to demonstrate and test the ability of ASAs powered by different LLMs to autonomously complete the entire research process.
-Each mission includes multiple sub-tasks, covering the full spectrum from conducting experiments to analyzing data and writing reports.
-The research plans (RPs) for these missions are shown in Figure 2.
-Among these RPs, RP 1 is the simplest involving only basic research steps for simulations given below.
+Among RP 1-3, RP 1 is the simplest, involving only basic research steps for simulations given below.
   • RP 1 requires generating a Python program to simulate a random walk, sampling different numbers of chain segments N, deriving the scaling relation $\left\langle R^{2}\right\rangle\propto N^{v}$, saving chain conformation graphs and scaling relation fit plots, and writing a research report.
-We established several evaluation criteria based on the requirements of the RP, such as whether graphs and plots were generated and whether the derived scaling relationships were correct, to assess task completion.
-We tested nine ASAs powered by different LLMs, including GPT-4o, conducting 20 trials for each RP and recording the achievement of the evaluation criteria across these trials, as shown in Figure 3.
 ASA-GPT-4o and ASA-GPT-4-Turbo excelled on RP 1, achieving near-perfect completion in both simulation and report writing tasks (Figure 3B).
 ASA-Claude-3.5 and ASA-Qwen-2 followed, performing well in simulation tasks but sometimes generating reports in response or as text files, failing to meet RP 1’s requirements.
-Other models performed well in plotting graphs but lacked accuracy in establishing correct physical models and deriving correct scaling values through simulations.
-Although chain conformation simulations are convenient for comparing different LLMs, they are less convincing due to being previously well-studied topics.
-To make a stronger case, we need more challenging problems.
-Additionally, to demonstrate that ASA can solve problems across different scientific domains, we designed RPs for gravitational simulation problems corresponding to RP S1 and RP S2 (Figure S2 and Figure 1C).
-The mission of RP S2 is to evaluate the risk posed by a meteorite (or asteroid) to earth, given its initial position and velocity and write a report.
-This mission is challenging even for graduates from physics departments.
-By simply inputting RP S2 into the same ASA used for chain conformation problems, the ASA will automatically acquire the necessary parameters of the solar system, code the simulation program to calculate the asteroid’s trajectory, and ultimately write the report.
-This result underscores the effectiveness of our API programming strategy (see Method section) and strongly demonstrates that ASA can address problems in entirely different fields.
+Other models performed well in plotting graphs but lacked accuracy in establishing the correct physical models and deriving correct scaling values through simulations.
 Automating Simulations with Remote Server Operations and Volume Exclusion
-RP 2 increases complexity with server interactions and RP 3 doubles the mission length and simulation difficulty. Brief introductions of these two RPs are shown below (Figure 2).
+RP 2 increases complexity with server interactions, and RP 3 doubles the mission length and simulation difficulty. Brief introductions of these two RPs are shown below (Figure 2).
   • RP 2 directly provides a random walk simulation program, asking the ASA to modify it, run simulations in a designated folder on a remote server, download the experimental data, and generate graphs and plots and a research report.
@@
 RP 3 required simulating a self-avoiding walk based on the random walk simulation in RP 1.
-Although the process was longer, it was less complex than RP 2, resulting in improved performance across agents.
+Although the process was longer, it was less complex than RP 2, resulting in an improved performance across agents.
 ASA-GPT-4o, ASA-GPT-4-Turbo, ASA-Claude-3.5, and ASA-Gemini-1.5-Pro performed relatively well.
@@
 Moreover, ASA-GPT-4o and ASA-Claude-3.5 demonstrated unique advantages in generating rich content.
-They frequently produced detailed reports exceeding 1,000 words, delving into the distinctions between RW and SA scaling, and occasionally included precise literature citations.
+They frequently produced detailed reports exceeding 1000 words, delving into the distinctions between RW and SA scaling, and occasionally included precise literature citations.
 ASA-Claude-3.5 referenced accurate self-avoiding walk scaling relationships validated through theoretical derivations and experimental verifications, accompanied by brief analyses of simulation result deviations.
 Additionally, ASA-GPT-4o presented experimental data in enriched formats, such as frequency distribution plots of end-to-end distances and interactive web-based charts.
-Automating Agent Coordination through RP Generation
-In the preceding section, RPs were provided by humans.
-In this section, we task a Main AI with automatically breaking down a human-provided RP into sub-tasks and distributing them as AI RPs to several Subordinate AIs.
-We provided RP 1 and 2 directly to the Master AI, adjusting the ASA to ensure each sub-task was handled by a distinct Subordinate AI without access to the Main AI’s or other Subordinate AIs’ conversation histories.
-The Main AI solely presents the AI RPs (Figure S1A) and received reports upon task completion.
-We also employed GPT-4-turbo for this experiment.
-The results (SI-data-2) revealed that for RP 1, the Main AI effectively decomposes the mission into sub-tasks and successfully distributed them to Subordinate AIs, thereby completing RP 1.
-However, for RP 2, the Main AI encounters repeated failures.
+Automating Agent Coordination through Sub-RP Generation
+In the preceding section, the ASA used a single AI to execute all of the content provided in the RP given by humans.
+In this section, we adjusted the logic for storing and managing dialogue history with the LLM within the ASA.
+We introduced a Manager AI that automatically breaks down a human-provided RP into subtasks and distributes them as sub-RPs to multiple Executor AIs.
+Each Executor AI is unaware of the dialogue history of the Manager AI and other Executor AIs, focusing solely on executing the assigned sub-RP.
+In contrast, the Manager AI solely presents the sub-RPs (Figure S1A) and receives reports upon task completion.
+We used GPT-4-Turbo for the experiments, testing ASA’s completion of RP 1 and RP 2 using the Manager-Executor AI mode.
+The results (SI-data-2) revealed that for RP 1, the Manager AI effectively decomposes the mission into subtasks and successfully distributes them to Executor AIs, thereby completing RP 1.
+However, for RP 2, the Manager AI encounters repeated failures.
 The primary issue lies in precise information transmission.
-RP 2 provides a random walk simulation program and details of a remote server, including hostname and username.
-Despite multiple attempts, the Main AI struggles to convey both the simulation program and server details accurately to the Subordinate AIs (Figure S1B).
-We attempted to instruct the Main AI to include all relevant information in the AI RPs, but this did not notably enhance the accuracy of information transmission.
-Additionally, we observed instances where AI RPs contained irrelevant information, potentially confusing the Subordinate AIs about the mission’s focus.
-Automating the Automation: Crafting Multi-tier RPs for Large-Scale Research
-Back in the Automating Basic Simulations section, for each ASA and each RP, we undertake 20 rounds of ‘automation’ by executing RP 1 and collecting the results including codes and files generated by the ASA.
-Can we automate the entire process of this study on reaserch automation or can we even automate the automation itself?
+RP 2 provides a random walk simulation program and details of a remote server, including the hostname and username.
+Despite multiple attempts, the Manager AI struggles to convey both the simulation program and server details accurately to the Executor AIs (Figure S1B).
+We attempted to instruct the Manager AI to include all relevant information in the sub-RPs, but this did not notably enhance the accuracy of the information transmission.
+Additionally, we observed instances where sub-RPs contained irrelevant information, potentially confusing the Executor AIs about the mission’s focus.
+Automating the Automation: Crafting Multitier RPs for the Large-Scale Research Process
+Back in the Automating Basic Simulations section, for each ASA and each RP, we undertake 20 rounds of “automation” by executing RP 1 and collecting the results, including codes and files generated by the ASA.
+Can we automate the entire process of this 20-round test on research automation, or can we automate the process of automation itself?
 RP 4 is designed for this purpose.
-It is nested within RP 1, instructing the Primary AI to execute the command “python AutoProg.py -s p1.txt -n i” 20 times.
-Each execution generates an Agent AI to fulfill RP 1 contained in p1.txt, while the Primary AI collects all generated files and conducts result analysis.
+It is nested within RP 1, instructing a Primary AI to execute the command “python AutoProg.py -s p1.txt -n” 20 times.
+Each execution generates a Secondary AI to fulfill RP 1 contained in p1.txt, while the Primary AI collects all generated files and conducts result analysis.
 The detailed content of RP 4 is depicted in Figure S1C.
@@
 The results (SI) strikingly demonstrate that the Primary AI successfully executed the 20 instances of RP 1, meticulously organizing all files into their respective folders and providing a brief summary after all experiments.
-During this process, the AI independently and without interruption wrote 66 programs (including 26 error versions), conducted more than 20 simulations, generated 120 images, and wrote 20 research reports totaling over 5,000 words.
-Marvelously, all these missions were accomplished without any human intervention.
-This illustrates the remarkable capability of multi-tier RP structures in handling large-scale missions.
+During this process, ASA-GPT-4-Turbo independently and without interruption wrote 66 programs (including 26 error versions), conducted more than 20 simulations, generated 120 images, and wrote 20 research reports totaling over 5000 words.
+Marvelously, all of these missions were accomplished without any human intervention.
+This illustrates the remarkable capability of multitier RP structures in handling large-scale missions.
+Promoting AI’s Autonomy—Brief-RP ASA
+To enhance the autonomy and flexibility of the AI, we developed the brief-RP ASA.
+We provide the system with a concise description of the research plan (RP) in the form of one or two sentences, allowing the AI to generate a detailed RP, including parameter settings, and subsequently execute the plan (see Section II in the SI for further details).
+Using brief-RP ASA, we addressed three distinct research questions (Figures S4-S6), such as predicting the viscosity of polymers based on a given data set using two machine learning methods and investigating the impact of natural disaster frequency on the dynamics of a simple grassland ecosystem.
+These examples illustrate that AI powered by large language models (LLMs) can exhibit a degree of creativity and autonomy in simulation-based research and that our ASA can be effectively applied to a wide range of research problems.
+While ASA was able to produce valuable insights in ecosystem studies, such as identifying a threshold above which increased disaster frequency leads to extinction events, its ability to generate entirely novel knowledge remains limited, and the reliability of its results requires further validation.
+Nevertheless, as LLM capabilities continue to advance, the potential of brief-RP ASA is boundless, enabling the rapid execution of a large number of simulation tasks.
 Discussion
-The experiments demonstrate that LLMs, with their robust coding ability 24, 32, 33, 34, can be directly employed in developing agents to autonomously complete complex research missions when aided by proper API programming.
-ASA-GPT-4o and ASA-GPT-4-turbo, notably, exhibits near-perfect completion rates for simpler missions, highlighting their potential.
-Multi-tier RP designs effectively handle more complex missions, though accurate information transmission remains challenging for such RPs.
+The experiments demonstrate that LLMs, with their robust coding ability,43, 51, 52, 53 can be directly employed in developing agents to autonomously complete complex research missions when aided by proper prompt engineering.
+ASA-GPT-4o and ASA-GPT-4-turbo, notably, exhibit near-perfect completion rates for simpler missions, highlighting their potential.
+Multitier RP designs effectively handle more complex missions, though accurate information transmission remains challenging for such RPs.
 We emphasize that these simulation tasks, particularly those involving remote server execution, cannot be accomplished through interactions with web-based LLMs alone.
@@
 Recommended Practices for Fully Autonomous Research
-ASA achieves fully autonomous research through continuous interaction with the LLM according to a specific logical design guided by the research plan (RP) provided by the researchers.
-We have identified several key practices in ASA logic design and RP design to enhance execution success rates.
+ASA achieves an autonomous research process guided by human-provided RPs through continuous interaction with the LLM according to a specific logical design.
+We identified several key practices in ASA logic design and RP design to enhance execution success rates.
   • Structure agent responses for efficient code and task extraction.
-Require the LLM to produce structured outputs, enabling ASA to accurately extract Python code, subtasks, and progress indicators.
+Prompt the LLM to produce structured outputs, enabling ASA to accurately extract Python code, subtasks, and progress indicators.
 This improves code execution success rates and ensures smooth mission progression.
@@
 While the LLM is powerful in code generation, it may struggle with complex programs on the first attempt.
-By integrating self-review and debugging process, the LLM can generate correct and task-compliant code within several dialogue rounds.
+By integrating self-review and debugging processes, ASA can generate correct and task-compliant code within several dialogue rounds.
   • Ensure key outputs at each subtask step for continuity.
@@
 Providing sufficient information in RP can improve mission success rates.
-In addition to necessary numeric/textual information, sometimes offering specific methods can guide the LLM to write correct programs more efficiently.
+In addition to necessary numeric/textual information, sometimes offering specific methods can guide ASA to write correct programs more efficiently.
 In summary, while the LLM has strong generative capabilities, its unpredictable outputs can complicate automation.
-ASA, based on Python’s AutoProg, though less intelligent, is reliable when logic is well-designed.
-By effectively structuring logic, ASA can oversee the LLM’s output, enhancing process reliability and achieving fully autonomous research.
+ASA, through prompt engineering to normalize LLM responses, although it restricts the LLM’s freedom, is reliable when logic is well-designed.
+By effectively structuring logic, ASA can oversee the LLM’s output, enhancing the smoothness and reliability of the automated simulation research process (Figure 4).
 Fig.
-4: Recommended practices for fully autonomous research.
+4: Recommended practices for automated simulation research workflow.
 By conducting iterative dialogues with the LLM adhering to a specific design logic, the success rate and efficiency of the automated process are enhanced.
@@
 Common Issues in Sequential Task
-Despite varying completion rates among tested ASAs, their errors and failure points share common characteristics, reflecting the inherent logical features of LLMs.
+Despite varying completion rates among tested ASAs, their errors and failure points share common characteristics, reflecting the inherent logical features of the LLMs.
 First, some ASAs tend to overlook parts of the RP’s requirements or “take shortcuts”, not fully following the RP’s instructions.
@@
 These issues affect mission completeness and can disrupt the entire workflow.
-To improve autonomy in AI for science, we need to address these issues by continuously generating outputs describing the current mission status, periodically reviewing generated data, and confirming strict adherence to mission requirements.
-These improvements can be integrated through AutoProg modifications or by incorporating them into the "system" prompt input to LLMs or fine-tuning the LLMs.
+To improve the smoothness of the automated process, we need to address these issues by continuously generating outputs describing the current mission status, periodically reviewing generated data, and confirming strict adherence to mission requirements.
+These improvements can be integrated through prompt engineering modifications or by incorporating them into the “system” prompt input to LLMs or fine-tuning the LLMs.
 Second, ASAs often underperform on technical aspects of tasks due to a lack of specialized domain knowledge.
@@
 In another example, for simulating self-avoiding walks, ASAs often incorrectly extend the random walk strategy by merely checking distances between points, resulting in high-energy state samples with large statistical errors.
-Correct sampling should use importance sampling to generate statistically significant conformations35.
-To avoid these issues in research, reliable solutions should be provided in the RPs, or relevant domain knowledge should be used for fine-tuning the LLMs to enhance their understanding.
+Correct sampling should use importance sampling to generate statistically significant conformations.54 To avoid these issues in research, reliable solutions should be provided in the RPs, or relevant domain knowledge should be used for fine-tuning the LLMs to enhance their understanding.
 Error Loops in Debugging
-Each AutoProg involves a Python code debugging process.
-When bugs are detected, the LLM receives error messages and attempts to revise the code, iterating until it executes correctly.
-However, we’ve noted instances where the LLM struggles to generate functional code, resulting in mission failures.
+The prompt engineering logic design of ASA includes a Python code debugging process.
+When bugs are detected, error messages are passed to the LLM, which is then prompted to revise the code.
+This process is iterative until the code executes correctly.
+However, we have noted instances where ASA struggles to generate functional code, resulting in mission failures.
 This typically occurs when the LLM becomes “stuck” after repeated unsuccessful debugging attempts, cycling through the same erroneous code without progress.
 This phenomenon may be associated with the LLM’s propensity to depend heavily on past dialogue content.
-Research has documented instances where LLMs utilize context to deliver responses preferred by users, indicating an excessive reliance on previous information 36, 37, 38.
-While this capability allows LLMs to adapt responses based on short dialogue histories, it can also lead to persistent errors.
+Research has documented instances where LLMs utilize context to deliver responses preferred by users, indicating an excessive reliance on previous information.55, 56, 45 While this capability allows LLMs to adapt responses based on short dialogue histories, it can also lead to persistent errors.
 In our experiments, to prevent such deadlocks, we implemented a maximum debug attempt limit.
-Once this limit is reached, the AutoProg returns a “mission failed” message.
+Once this limit is reached, ASA’s AutoProg returns a “mission failed” message.
 To ensure the smooth execution of concrete research missions, it is crucial to avoid these debugging loops.
-One solution involves clearing parts of the memory after a certain number of debug attempts or encouraging the LLM to generate varied content in each iteration (Some LLMs include parameters to avoid generating duplicate content and to adjust the creativity of the models.).
+One solution involves clearing parts of the memory after a certain number of debug attempts or prompting the LLM to generate varied content in each iteration (Some LLMs include parameters to avoid generating duplicate content and to adjust the creativity of the models.).
 Balancing Global Oversight and Local Attention
-In our experiments, ASAs exhibited a significant level of global oversight, which allows them to track mission progress based on the dialogue history.
-They can retain details from earlier stages of missions even after focusing on lengthy sub-tasks.
-However, as discussed above, ASAs occasionally skip steps in prolonged missions or fail to deliver comprehensive solutions for sub-tasks, indicating a lack of sufficient local attention.
+In our experiments, ASA exhibited a significant level of global oversight, which allows it to track mission progress based on the dialogue history.
+ASA can retain details from earlier stages of missions even after focusing on lengthy subtasks.
+However, as discussed above, ASA occasionally skips steps in prolonged missions or fails to deliver comprehensive solutions for subtasks, indicating a lack of sufficient local attention.
 Achieving successful and precise mission completion requires a delicate balance between global oversight and local attention, two perspectives that are normally conflicting.
-In Automating Agent Coordination section, we designed a collaboration plan where a Main AI and some Subordinate AIs handled different aspects of the mission: the Main AI defined sub-task requirements and received reports, while Subordinate AIs focused on specific sub-tasks.
-This approach minimized redundant information in their dialogue history, enhancing the Main AI’s global awareness and the Subordinate AIs’ local attention.
-While the Main AI effectively completed the mission in RP 1, the Main-Sub AI model encountered challenges in RP 2 due to the failure of precise information transmission.
-One potential optimization is to involve human-designed RPs distributed among various Subordinate AIs for execution.
+In the Automating Agent Coordination through sub-RP Generation section, we designed a collaboration plan where a Manager AI and some Executor AIs handled different aspects of the mission: the Manager AI defined subtask requirements and received reports, while Executor AIs focused on specific subtasks.
+This approach minimized redundant information in different AIs’ dialogue history, enhancing the Manager AI’s global awareness and the Executor AIs’ local attention.
+While the Manager AI effectively completed the mission in RP 1, the Manager-Executor AI mode encountered challenges in RP 2 due to the failure of precise information transmission.
+One potential optimization involves human-designed RPs distributed among various Executor AIs for execution.
 Limitations and Next Steps
-This study demonstrates the capability of LLM powered research agent ASA to independently undertake sophisticated simulation research missions, using a well-recognized polymer physics simulation problem and a more challenging celestial-simulation problem as representative examples.
-However, there remains ample opportunity to explore ASA’s potential in tackling novel scientific challenges.
-Additionally, this study focuses on presenting an automated workflow, leaving room for further optimization of prompts and automation programs.
-Several inherent limitations of ASAs were identified: (1) ASAs may overlook mission requirements, necessitating methods to prioritize critical information; (2) general LLMs lack specialized domain knowledge, requiring supplementary information in RPs; (3) LLMs may persist in errors by relying too heavily on previous content, requiring memory management strategies in ASAs’ development; (4) balancing global awareness and local attention poses a challenge; (5) accurately transmitting specific information remains problematic.
-Addressing these issues is crucial for enhancing the role of AI in scientific research.
-This requires a thorough analysis of these limitations and adjustments to research methodologies and LLM capabilities, leveraging their strengths while mitigating their weaknesses.
-Conclusion
-This study demonstrates the potential of LLM powered research agent ASA to autonomously conduct scientific research missions, using a polymer theory simulation problem and a more challenging asteroid-trajectory prediction problem as model problems.
-This process, encompassing simulation coding, remote execution, data analysis, and report generation, represents the first fully autonomous AI-directed simulation research: once the research plan is made, only a single command (e.g., ‘do prj.txt’) is required to initiate and complete the process.
-The notable success herein can be largely attributed to the sophisticated API programming.
-Agents such as ASA-GPT-4o demonstrate high completion rates for missions, a strong grasp of task requirements, and consistent alignment with overall objectives across extensive tasks.
-Notably, the performance exhibited in completing RP 4, entailing a hierarchy of nested tasks, surpassed expectations in a significant manner.
-These capabilities underscore the readiness of ASAs to independently manage long-term scientific research missions, offering a new perspective on AI applications in science as well as in chemical engineering.
+This study demonstrates the capability of an autonomous simulation agent (ASA), constructed through LLM prompt engineering and an automated program, to automatically complete long-task workflows according to a Research Plan (RP) provided by humans.
+We used a well-recognized polymer physics simulation problem and a more challenging celestial simulation problem as representative examples.
+We tested different LLMs to construct ASAs, and while agents like ASA-GPT-4o, ASA-GPT-4-turbo, and ASA-Claude-3.5 demonstrated very high completion rates in the test tasks, we also identified several inherent limitations of the ASA method in handling long tasks: (1) Mission Requirement Overlooking.
+ASAs may overlook mission requirements, necessitating methods to prioritize critical information; (2) Lack of Specialized Domain Knowledge.
+General LLMs lack specialized domain knowledge, requiring supplementary information in RPs; (3) Error Persistence.
+LLMs may persist in errors by relying too heavily on previous content, requiring memory management strategies in ASAs’ development; (4) Balancing Global Awareness and Local Attention.
+Ensuring that ASA maintains a broad overview while focusing on specific details as needed poses a challenge; and (5) Information Transmission Accuracy.
+Securing that specific information is correctly interpreted and utilized remains problematic.
+These limitations highlight areas for further improvement and research to enhance the robustness and reliability of ASAs in handling complex, end-to-end simulation tasks.
+Despite the current limitations of LLMs, which mean that the scientific conclusions drawn by ASA are not yet fully reliable, these results showcase the applicability of the ASA approach across various fields (Figures S4-S7) and its potential to assist in exploring new questions.
+Further, we need to endow AI with more autonomous judgment and exploration capabilities, enabling it to summarize existing information and continuously generate new RPs for execution in more complex research scenarios, so as to tackle novel scientific challenges in the future.
+Conclusions
+This study demonstrates the potential of LLM-powered research agent ASA to autonomously conduct end-to-end scientific research missions under the guidance of an RP provided by humans.
+This comprehensive simulation research process, encompassing simulation coding, remote execution, data analysis, and report generation, represents an autonomous AI-conducted simulation research process.
+Once the research plan is made, only a single command (e.g., “do prj.txt”) is required to initiate and complete the process.
+The success herein can be largely attributed to prompt engineering and automated program design.
+We tested ASAs equipped with different LLMs using both a well-known polymer chain simulation problem and a more challenging celestial simulation problem.
+Human-written RPs were provided, requiring the ASAs to execute according to their content, and the task completion of each agent was evaluated.
+Agents such as ASA-GPT-4o demonstrated high completion rates, a strong grasp of task requirements, and consistent alignment with the overall objectives across extensive tasks.
+Notably, the performance exhibited in completing RP 4, which entailed a hierarchy of nested tasks, surpassed expectations in a significant manner.
+These capabilities underscore the readiness of ASA to conduct long-term research missions, offering a new perspective on AI applications in science, particularly in simulation studies, to enhance research efficiency.
 Methods
 API Automation Program (AutoProg)
-Based on the official documentation of various LLMs, we implemented the API automation programs (AutoProg) by configuring Python API calling methods and setting up the local environment.
-This setup included installing necessary Python packages like numpy, matplotlib, paramiko, and python-docx tailored for our simulation tasks.
-An AutoProg serves dual functions: 1) it engages in dialogue with the LLM via the API and preserves all historical dialogues, and 2) it extracts and executes Python programs generated by the LLM in response to prompts.
-For instance, in the case of RP 1, an AutoProg initiates the automated research process by reading mission requirements saved in a txt file.
-It then calls the API to transmit the RP to the LLM, instructing it to generate a comprehensive Python program.
-The AutoProg extracts this program from the LLM’s response, sends it back to the LLM via the API along with entire dialogue history, and requests error checking and compliance verification against mission requirements.
-Subsequently, the AutoProg executes the validated Python program, capturing and storing output or errors.
-In case of errors, the AutoProg sends the prior dialogues and program errors back to the LLM, asking for modifications.
-This process repeats until the program runs correctly.
+The ASA leverages an API automation program (AutoProg) to enable multiround text interactions with LLMs via API (Figure 1) and to execute the programs written by the LLM locally.
+AutoProg is designed with a logical prompt flow that guides the LLM to respond with specific content at different stages based on the conversation history and prompts, including prompting the LLM to break down the RP into subtasks, generate Python programs, debug the programs, and check the completion status of the tasks.
+For instance, in the case of RP 1, ASA’s AutoProg initiates the automated simulation process by reading the RP.txt file written by a human.
+It then calls the API to transmit the content of the RP to the LLM, while prompting it to break the plan down into multiple subtasks and write Python code to complete the first subtask in its response.
+The AutoProg extracts the Python program from the LLM’s response, sends it back to the LLM via the API along with entire dialogue history, and requests error checking and compliance verification against mission requirements.
+Subsequently, AutoProg executes the validated Python program, capturing and storing output or errors.
+In the case of errors, the AutoProg sends the prior dialogues and program errors back to the LLM, asking for modifications.
+This process is repeated until the program runs correctly.
 If errors persist beyond a predefined threshold, the AutoProg halts with a “mission failed” output.
 Conversely, upon successful execution, it returns the program output to the LLM, prompting it to check the dialogues and files in the working directory to determine the task progress.
-If the task remains incomplete, the next Python program will be generated by the LLM to continue the mission, iterating through checking and debugging processes until completion.
-When the mission achieves full execution, the AutoProg concludes with a "mission complete" output.
-AutoProg.py and RP.txt, containing predefined instructions, are stored in the local working directory.
-To initiate the automated research process for each experiment, users navigate to the directory via the command line (CMD) and execute commands such as “python AutoProg.py -s RP.txt”.
-Alternatively, AutoProg can be elevated to a command by creating a batch file (Windows) or shell script (Linux), allowing users to simply enter “AutoProg RP.txt” from any directory to execute it.
-A detailed video recording of the execution process is available for reference (SI-video), in which a batch file is used to facilitate the execution by running the command “do AI4S_prj1.txt”.
-To evaluate LLM performance across RP 1-3, each ASA underwent 20 tests per RP, with all generated files and AutoProg outputs saved. Sample results are provided in the data (SI-data-1).
+If the task remains incomplete, it prompts the LLM to generate the next Python program to continue the mission, iterating through checking and debugging processes until completion.
+When the mission achieves full execution, the AutoProg concludes with a “mission complete” output.
+AutoProg.py and RP.txt are stored in the local working directory.
+To initiate the automated simulation process using ASA for each experiment, users navigate to the directory via the command line (CMD) and execute commands such as “python AutoProg.py -s RP.txt”.
+Alternatively, AutoProg can be elevated to a command by creating a batch file (Windows) or shell script (Linux), allowing users to simply enter “command RP.txt” from any directory to execute it.
+A detailed video recording of the execution process is available for reference (SI-Video), in which a batch file is used to facilitate execution by running the command “do AI4S_prj1.txt”.
+To evaluate the ASA performance powered by different LLMs across RP 1-3, each ASA underwent 20 tests per RP, with all generated files and outputs saved.
+Sample results are provided in the data (SI-data-1).
+The ASA we developed is a general system (minor adjustments may be required in AutoProg for different LLM APIs) that requires only a change in the original RP to be applicable to any simulation problem involving Python programming.
+In addition to the RPs designed for polymer conformation simulations, we have also crafted RPs for solar system planetary orbit simulations and asteroid trajectory simulations (Figure 1C and SI).
 ASA Scoring via EWM and TOPSIS
@@
 For quantitative analysis and scoring, we deployed the Entropy Weight Method (EWM) and TOPSIS (Technique for Order of Preference by Similarity to Ideal Solution).
-The EWM, a well-established approach for determining the relative significance of criteria in decision-making 39, was utilized to assign weights based on data variability.
+The EWM, a well-established approach for determining the relative significance of criteria in decision-making, 57 was utilized to assign weights based on data variability.
 This ensures criteria with higher variability, and thus more informational value, receive higher weights in the final evaluation.
-TOPSIS, originally formulated by Hwang and Yoon 40, was subsequently applied.
+TOPSIS, originally formulated by Hwang and Yoon,58 was subsequently applied.
 This method computes the geometric distance between each agent’s performance and both ideal and anti-ideal solutions, facilitating a comparative assessment of ASA efficacy.
@@
 Initially, fulfillment data for all ASAs across various criteria were gathered over 20 trials, encapsulated in matrix X, wherein $x_{ij}$ signifies the frequency of criterion j’s attainment by agent i.
-Normalization was then performed on $x_{ij}$:
+Normalization was then performed on $x_{ij}$
 $r_{ij}=\frac{x_{ij}-min(x_{ij})}{max(x_{ij})-min(x_{ij})},$ (1)
-yielding the normalized matrix $\mathbf{R}=[r_{ij}]$. Subsequently, criterion weights were established utilizing EWM. The relative proportion $p_{ij}$ of the criterion j:
+yielding the normalized matrix $\mathbf{R}=[r_{ij}]$. Subsequently, criterion weights were established utilizing EWM. The relative proportion $p_{ij}$ of criterion j
 $p_{ij}=\frac{r_{ij}}{\sum_{i=1}^{m}r_{ij}}.$ (2)
-The entropy $e_{j}$ of the criterion j:
+The entropy $e_{j}$ of the criterion j
 $e_{j}=-k\sum_{i=1}^{m}p_{ij}\ln(p_{ij}),$ (3)
-with $k=1/\ln(m)$, and $m$ denoting the total number of agents.
-The weight $\omega_{j}$ of the criterion $j$:
+with $k=1/\ln(m)$ and $m$ denoting the total number of agents.
+The weight $\omega_{j}$ of the criterion $j$
 $\omega_{j}=\frac{d_{j}}{\sum_{j=1}^{n}d_{j}},$ (4)
@@
 The positive ideal solution $v_{j}^{+}$ and negative ideal solution $v_{j}^{-}$ were determined as $v_{j}^{+}=max(v_{ij})$ and $v_{j}^{-}=min(v_{ij})$, respectively.
-Then the geometric distances $S_{j}^{+}$ and $S_{j}^{-}$ from $v_{ij}$ to the positive and negative ideal solutions can be computed, respectively, as
+Then the geometric distances $S_{j}^{+}$ and $S_{j}^{-}$ from $v_{ij}$ to the positive and negative ideal solutions can be computed, respectively, as follows
 $S_{j}^{+}=\sqrt{\sum_{j=1}^{n}\left(v_{ij}-v_{j}^{+}\right)^{2}}$ (5)
@@
 $S_{j}^{-}=\sqrt{\sum_{j=1}^{n}\left(v_{ij}-v_{j}^{-}\right)^{2}}.$ (6)
-The relative closeness $C_{i}^{*}$ for each agent was calculated as:
+The relative closeness $C_{i}^{*}$ for each agent was calculated as follows
 $C_{i}^{*}=\frac{S_{j}^{-}}{S_{j}^{+}+S_{j}^{-}}$ (7)
@@
 ASSOCIATED CONTENT
+Data Availability Statement
+The source code and example results of ASA can be found at https://github.com/zokaraa/autonomous_simulation_agent.
 Supporting Information
-The following Supporting Information is available free of charge.
-Partial results for gravitational simulation mission along with examples of reports written by ASAs (SI.pdf); partial AutoProgs and experimental result files for RP 1-3, main AI generating RP, and Multi-tier human RP (Some information has been obscured; readers can replace it with their own information and follow the instructions to run the AutoProgs) (SI-data-1.zip and SI-data-2.zip); a demo video demonstrating the operation process of the automatic research system (SI-video.mp4).
+The following Supporting Information is available free of charge at https://pubs.acs.org/doi/10.1021/acs.jcim.4c01653.
+  • Supporting diagram for Manager-Executor AI mode and Multitier AI mode; partial results for gravitational simulation mission; and results for AI-designed RP along with examples of reports written by ASAs (PDF)
+  • A demo video demonstrating the operation process of the automatic research system (MP4)
+  • ASA’s AutoProg examples and experimental result files for RP 1-3 (ZIP)
+  • Manager-Executor AI mode, Multitier AI mode, and AI-designed RP (ZIP)
 AUTHOR INFORMATION
@@
 Author Contributions
-Z.H.L. was the primary author of the paper. Y.B.C. was responsible for part of the data analysis and statistics. J.F.L. originally initiated and supervised the project.
+Z.L. was the primary author of the paper. Y.C. was responsible for part of the data analysis and statistics. J.L. originally initiated and supervised the project.
 Notes
@@
 ACKNOWLEDGEMENTS
-J.F.L. acknowledge supports from National Natural Science Foundation of China (Nos. 52394272, 22373022) and National Key Research and Development Program of China (No. 2023YFA0915300).
+J.F.L.
+acknowledge supports from National Natural Science Foundation of China (Nos.
+52394272, 22373022) , National Key Research and Development Program of China (No.
+2023YFA0915300) and Shanghai Science and Technology Innovation Action Plan (No.
+24JD1400700).
 Notes and references
+  • Taylor et al. 2022 Taylor, R.; Kardas, M.; Cucurull, G.; et al. Galactica: A Large Language Model for Science. arXiv:2211.09085, 2022.
+  • Evans et al. 2021 Evans, R.; O’Neill, M.; Pritzel, A.; et al. Protein complex prediction with AlphaFold-Multimer. bioRxiv:2021.10.04.463034, 2021.
+  • Abramson et al. 2024 Abramson, J.; Adler, J.; Dunger, J.; et al. Accurate structure prediction of biomolecular interactions with AlphaFold 3. Nature 2024, 630, 493–500.
+  • Hu et al. 2024 Hu, J.; Zhou, L.; Jiang, J. Efficient Machine Learning Force Field for Large-Scale Molecular Simulations of Organic Systems. CCS Chemistry 2024, 1–15, Just Published.
+  • Xu and Jiang 2024 Xu, L.; Jiang, J.
+Synergistic Integration of Physical Embedding and Machine Learning Enabling Precise and Reliable Force Field.
+Journal of Chemical Theory and Computation 2024, 20, 7785–7795.
+  • 6 Li, J.; Zhang, H.; Chen, J. Z. Y. Structural Prediction and Inverse Design by a Strongly Correlated Neural Network. Physical Review Letters 123, 108002.
+  • Liu et al.
+2023 Liu, Z.; Liu, Y.-X.; Yang, Y.; Li, J.
+Template Design for Complex Block Copolymer Patterns Using a Machine Learning Method.
+ACS Applied Materials & Interfaces 2023, 15, 31049–31056.
+  • Qiu and Sun 2024 Qiu, H.; Sun, Z. On-demand reverse design of polymers with PolyTAO. npj Computational Materials 2024, 10, 273.
+  • Qiu et al.
+2024 Qiu, H.; Wang, J.; Qiu, X.; Dai, X.; Sun, Z.-Y.
+Heat-Resistant Polymer Discovery by Utilizing Interpretable Graph Neural Network with Small Data.
+Macromolecules 2024, 57, 3515–3528.
+  • Himanen et al. 2019 Himanen, L.; Geurts, A.; Foster, A. S.; Rinke, P. Data-Driven Materials Science: Status, Challenges, and Perspectives. Advanced Science 2019, 6, 1900808.
+  • Wang et al. 2023 Wang, H.; Fu, T.; Du, Y.; et al. Scientific discovery in the age of artificial intelligence. Nature 2023, 620, 47–60.
+  • OpenAI 2023 OpenAI GPT-4 Technical Report. https://cdn.openai.com/papers/gpt-4.pdf, 2023; Retrieved June 24, 2024.
+  • Anthropic 2023 Anthropic Model Card and Evaluations for Claude Models.
+https://www-cdn.anthropic.com/files/4zrzovbb/website/bd2a28d2535bfb0494cc8e2a3bf135d2e7523226.pdf, 2023; Retrieved June 24, 2024.
+  • Xu et al.
+2022 Xu, F.
+F.; Alon, U.; Neubig, G.; Hellendoorn, V.
+J.
+A systematic evaluation of large language models of code.
+Proceedings of the 6th ACM SIGPLAN International Symposium on Machine Programming.
+New York, NY, USA, 2022; pp 1–10.
+  • Hou et al.
+2024 Hou, X.; Zhao, Y.; Liu, Y.; Yang, Z.; Wang, K.; Li, L.; Luo, X.; Lo, D.; Grundy, J.; Wang, H.
+Large Language Models for Software Engineering: A Systematic Literature Review.
+ACM Transactions on Software Engineering and Methodology 2024,
   • Devlin et al.
@@
   • Thoppilan et al. 2022 Thoppilan, R. et al. LaMDA: Language Models for Dialog Applications. arXiv:2201.08239, 2022.
-  • OpenAI 2023 OpenAI GPT-4 Technical Report. https://cdn.openai.com/papers/gpt-4.pdf, 2023; Retrieved June 24, 2024.
-  • Anthropic 2023 Anthropic Model Card and Evaluations for Claude Models.
-https://www-cdn.anthropic.com/files/4zrzovbb/website/bd2a28d2535bfb0494cc8e2a3bf135d2e7523226.pdf, 2023; Retrieved June 24, 2024.
   • Ramesh et al.
@@
   • Kasneci et al. 2023 Kasneci, E. et al. ChatGPT for good? On opportunities and challenges of large language models for education. Learning and Individual Differences 2023, 103, 102274.
-  • Himanen et al. 2019 Himanen, L.; Geurts, A.; Foster, A. S.; Rinke, P. Data-Driven Materials Science: Status, Challenges, and Perspectives. Advanced Science 2019, 6, 1900808.
-  • Taylor et al. 2022 Taylor, R.; Kardas, M.; Cucurull, G.; et al. Galactica: A Large Language Model for Science. arXiv:2211.09085, 2022.
-  • Wang et al. 2023 Wang, H.; Fu, T.; Du, Y.; et al. Scientific discovery in the age of artificial intelligence. Nature 2023, 620, 47–60.
-  • Evans et al. 2021 Evans, R.; O’Neill, M.; Pritzel, A.; et al. Protein complex prediction with AlphaFold-Multimer. bioRxiv:2021.10.04.463034, 2021.
-  • Abramson et al. 2024 Abramson, J.; Adler, J.; Dunger, J.; et al. Accurate structure prediction of biomolecular interactions with AlphaFold 3. Nature 2024, 630, 493–500.
   • Brown et al. 2019 Brown, N.; Fiscato, M.; Segler, M.; et al. GuacaMol: Benchmarking Models for de Novo Molecular Design. Journal of Chemical Information and Modeling 2019, 59, 1096–1108.
@@
 Nature Communications 2024, 15, 4705.
+  • Guo et al.
+2024 Guo, E.; Gupta, M.; Deng, J.; Park, Y.-J.; Paget, M.; Naugler, C.
+Automated Paper Screening for Clinical Reviews Using Large Language Models: Data Analysis Study.
+Journal of Medical Internet Research 2024, 26, e48996.
+  • Bolaños et al. 2024 Bolaños, F.; Salatino, A.; Osborne, F.; others Artificial intelligence for literature reviews: opportunities and challenges. Artificial Intelligence Review 2024, 57, 259.
+  • Lozano et al. 2024 Lozano, A.; Fleming, S. L.; Chiang, C.-C.; Shah, N. Biocomputing 2024; 2024; pp 8–23.
+  • Dagdelen et al. 2024 Dagdelen, J.; Dunn, A.; Lee, S.; others Structured information extraction from scientific text with large language models. Nature Communications 2024, 15, 1418.
+  • Fink et al.
+2023 Fink, M.
+A.; Bischoff, A.; Fink, C.
+A.; Moll, M.; Kroschke, J.; Dulz, L.; Heußel, C.
+P.; Kauczor, H.-U.; Weber, T.
+F.
+Potential of ChatGPT and GPT-4 for Data Mining of Free-Text CT Reports on Lung Cancer.
+Radiology 2023, 308, e231362.
+  • Ma et al.
+2023 Ma, P.; Ding, R.; Wang, S.; Han, S.; Zhang, D.
+InsightPilot: An LLM-Empowered Automated Data Exploration System.
+Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing: System Demonstrations.
+Singapore, 2023; pp 346–352.
+  • Peng et al. 2024 Peng, D.; Zheng, L.; Liu, D.; others Large-language models facilitate discovery of the molecular signatures regulating sleep and activity. Nature Communications 2024, 15, 3685.
+  • Li et al. 2024 Li, T.; Shetty, S.; Kamath, A.; others CancerGPT for few shot drug pair synergy prediction using large pretrained language models. npj Digital Medicine 2024, 7, 40.
+  • Subba et al.
+2024 Subba, B.; Toufiq, M.; Omi, F.; others Human-augmented large language model-driven selection of glutathione peroxidase 4 as a candidate blood transcriptional biomarker for circulating erythroid cells.
+Scientific Reports 2024, 14, 23225.
+  • Liu et al. 2023 Liu, F.; Zhu, T.; Wu, X.; others A medical multimodal large language model for future pandemics. npj Digital Medicine 2023, 6, 226.
+  • Madani et al. 2023 Madani, A.; Krause, B.; Greene, E. R.; others Large language models generate functional protein sequences across diverse families. Nature Biotechnology 2023, 41, 1099–1106.
   • Jablonka et al. 2024 Jablonka, K. M.; Schwaller, P.; Ortega-Guerrero, A.; et al. Leveraging large language models for predictive chemistry. Nature Machine Intelligence 2024, 6, 161–169.
@@
   • Zhou et al. 2024 Zhou, Z.; Ning, X.; Hong, K.; et al. A Survey on Efficient Inference for Large Language Models. arXiv:2404.14294, 2024.
-  • Abolhasani and Kumacheva 2023 Abolhasani, M.; Kumacheva, E. The rise of self-driving labs in chemical and materials sciences. Nature Synthesis 2023, 2, 483–492.
-  • Szymanski et al. 2023 Szymanski, N. J.; Rendy, B.; Fei, Y.; et al. An autonomous laboratory for the accelerated synthesis of novel materials. Nature 2023, 624, 86–91.
-  • Volk et al.
-2023 Volk, A.
-A.; Epps, R.
-W.; Yonemoto, D.
-T.; et al.
-AlphaFlow: autonomous discovery and optimization of multi-step chemistry using a self-driven fluidic lab guided by reinforcement learning.
-Nature Communications 2023, 14, 1403.
+  • Yao et al.
+2023 Yao, S.; Yu, D.; Zhao, J.; Shafran, I.; Griffiths, T.; Cao, Y.; Narasimhan, K.
+Tree of Thoughts: Deliberate Problem Solving with Large Language Models.
+Advances in Neural Information Processing Systems.
+2023; pp 11809–11822.
+  • Wei et al.
+2022 Wei, J.; Bosma, M.; Zhao, V.
+Y.; Guu, K.; Yu, A.
+W.; Lester, B.; Du, N.; Dai, A.
+M.; Le, Q.
+V.
+Finetuned Language Models are Zero-Shot Learners.
+The Tenth International Conference on Learning Representations, ICLR 2022, Virtual Event, April 25-29, 2022.
+2022.
+  • Kojima et al. 2022 Kojima, T.; Gu, S. S.; Reid, M.; Matsuo, Y.; Iwasawa, Y. Large Language Models are Zero-Shot Reasoners. Advances in Neural Information Processing Systems. 2022; pp 22199–22213.
   • Flory 1969 Flory, P. J. Statistical Mechanics of Chain Molecules; Interscience Publishers, 1969.
@@
   • Gao et al. 2020 Gao, T.; Fisch, A.; Chen, D. Making Pre-trained Language Models Better Few-shot Learners. arXiv:2012.15723, 2020.
-  • Wei et al.
-2022 Wei, J.; Bosma, M.; Zhao, V.
-Y.; Guu, K.; Yu, A.
-W.; Lester, B.; Du, N.; Dai, A.
-M.; Le, Q.
-V.
-Finetuned Language Models are Zero-Shot Learners.
-The Tenth International Conference on Learning Representations, ICLR 2022, Virtual Event, April 25-29, 2022.
-2022.
   • Ravindran et al. 1987 Ravindran, A.; Phillips, D. T.; Solberg, J. J. Operations Research Principles & Practice, 2nd ed.; John Wiley: New York, 1987.
   • Hwang and Yoon 1981 Hwang, C. L.; Yoon, K. Multiple Attribute Decision Making; Lecture Notes in Economics and Mathematical Systems; Springer: Berlin, Heidelberg, 1981; Vol. 186.
+For Table of Contents Use Only.
+By submitting a research plan file using the command do prj1.txt, ASA will automatically execute all subsequent simulation tasks without any human intervention.